### v1

# Understanding Outer Optimizers in Local SGD: Learning Rates, Momentum, and Acceleration

Abstract
Modern machine learning often requires training with large batch size, distributed data, and massively parallel compute hardware (like mobile and other edge devices or distributed data centers). Communication becomes a major bottleneck in such settings but methods like Local Stochastic Gradient Descent (Local SGD) show great promise in reducing this additional communication overhead. Local SGD consists of three parts: a local optimization process, an aggregation mechanism, and an outer optimizer that uses the aggregated updates from the nodes to produce a new model. While there exists an extensive literature on understanding the impact of hyperparameters in the local optimization process, the choice of outer optimizer and its hyperparameters is less clear. We study the role of the outer optimizer in Local SGD, and prove new convergence guarantees for the algorithm. In particular, we show that tuning the outer learning rate allows us to (a) trade off between optimization error and stochastic gradient noise variance, and (b) make up for ill-tuning of the inner learning rate. Our theory suggests that the outer learning rate should sometimes be set to values greater than $1$. We extend our results to settings where we use momentum in the outer optimizer, and we show a similar role for the momentum-adjusted outer learning rate. We also study acceleration in the outer optimizer and show that it improves the convergence rate as a function of the number of communication rounds, improving upon the convergence rate of prior algorithms that apply acceleration locally. Finally, we also introduce a novel data-dependent analysis of Local SGD that yields further insights on outer learning rate tuning. We conduct comprehensive experiments with standard language models and various outer optimizers to validate our theory.
1 Introduction
Training very large scale machine learning models requires a lot of compute. This compute is often centrally controlled by a single entity and tightly connected in a data center. Gradients are constantly synchronized, hardware failures are controlled and mitigated, and things (mostly) run smoothly. Building this training infrastructure is expensive, however, and centralized control might not be desirable for all models. This has led to a surge of interest in decentralized collaborative training of large-scale models across different, potentially poorly connected clusters [10, 23, 22]. This has motivated the adoption of federated learning algorithms in training language models, chiefly for scalability and communication efficiency rather than data privacy. Efficient parallelization strategies also factored in the remarkable recent training of DeepSeek V3 and R1 on a tight budget [7, 6].
A foundational algorithm in distributed and federated optimization is Local SGD [54]. Many popular algorithms fit in the FedOpt template [48] (Algorithm 1), including FedAdam [48], FedRR [40, 38], DiLoCo [10, 23] and many others. FedOpt solves the minimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ given access to $M$ different computational nodes and unbiased stochastic gradients of $f$. FedOpt consists of three main components: an inner update loop on every client, an aggregation of the client updates, and then an outer update step taken on the server.
1:Input. Update rules $\mathrm{LocalUpdate}$ and $\mathrm{OuterUpdate}$. Initial point $x_{0}$.
2:for communication rounds $r=0,1,\ldots,R-1$ do
3:  Broadcast $x_{r}$ to each node $m$
4:  for each node $m$ in parallel do
5:   Set $y_{m,r,0}=x_{r}$.
6:   for local steps $h=0,1,\ldots,H-1$ do
7:    Set $y_{m,r,h+1}=\mathrm{LocalUpdate}(y_{m,r,h},g_{m,r,h})$ for stochastic gradient $g_{m,r,h}$ at $y_{m,r,h}$.
8:   end for
9:   Communicate $y_{m,r,H}$ to the server.
10:  end for
11:  Compute the update or “outer gradient” $\hat{\Delta}_{r,H}=\frac{1}{M}\sum_{m=1}^{M}(y_{m,r,H}-x_{r})$.
12:  Update $x_{r+1}=\text{OuterUpdate}(x_{r},-\hat{\Delta}_{r,H})$.
13:end for
Algorithm 1 The FedOpt Algorithmic Template
When both the local and outer update rules correspond to gradient descent (i.e. $x_{\mathrm{new}}=x_{\mathrm{old}}-\beta\Delta$ for some stepsize $\beta$ and update vector $\Delta$), the corresponding algorithm is Generalized Local SGD. If we additionally take the outer stepsize to be $1$, we get Local SGD. Local SGD simply does $H$ steps of SGD on each node, and then averages the result after applying the updates. This is the most common form in which the algorithm is analyzed, as in e.g. [52, 27, 58, 28, 13, 46]. In practice, different choices of outer optimizers perform better. For example, DiLoCo/OpenDiLoCo use SGD with Nesterov Momentum as the outer optimizer [10]. This has motivated much analysis of different outer optimizers and their impact [48, 37, 24, 53]. However, our theoretical understanding of the fundamental Generalized Local SGD algorithm remains limited. In particular, it is not clear why the bilevel optimization structure of the algorithm is helpful from an optimization perspective, even in the i.i.d. setting where the data distribution is the same on all the nodes. Additionally and to the best of our knowledge, we have no explicit expressions for what the ideal learning rate pair $(\eta,\gamma)$ should be. Empirically, outer optimizers employing Nesterov acceleration have the best performance, yet to the best of our knowledge why or how it improves convergence is not known.
Contributions. Our paper takes steps to address the above questions and makes the following contributions.
  • We conduct a novel, tighter analysis of Generalized Local SGD (Theorem 3.3) that shows the outer learning rate plays a dual role. It (a) interpolates between two extreme regimes: taking many effective steps at the cost of higher variance to taking fewer steps but at reduced variance and (b) increases the algorithmic robustness to hyperparameter tuning by making up for ill-tuned inner learning rates. The latter holds even in the absence of any stochastic gradient noise.
  • We extend the above analysis to cover Generalized Local SGD where the outer optimizer also uses momentum (Theorem 3.5) and show that this gives additional leeway in tuning $\gamma$.
  • We provide a convergence analysis for Local SGD with an accelerated outer optimizer and unaccelerated inner optimizer (Theorem 3.6), showing that using Nesterov acceleration in the outer loop achieves better dependence on the number of communication rounds $R$ in the drift terms compared to standard Local SGD and improving upon the convergence rate of FedAc [60].
  • We also derive an adaptive, data-dependent, high-probability guarantee for the convergence of Local SGD with GD as the outer optimizer (Theorem 3.8) that shows further benefits of tuning the outer stepsize in more nuanced settings.
  • We additionally conduct an extensive empirical analysis for training large-scale language models with various outer optimizers (gradient descent, accelerated gradient descent, and Schedule-Free gradient descent).
We now review related work, then proceed to our main results.
2 Related Work
There is a rich literature on algorithms for communication-efficient distributed optimization for federated learning [29], where multiple clients collaborate on solving a machine learning problem  [54]. Federated learning algorithms are designed to reduce the effect of data heterogeneity [26, 54, 42], ensure the data stays private [56], deal with intermittent or cyclic client availability [12], among other issues.
[…]
The most popular algorithm in the federated optimization literature is Local SGD or Federated Averaging [54]. It is a generalization of minibatch SGD that, rather than communicating at every step of the optimization process, communicates only intermittently. Local SGD shows remarkable efficiency in many settings in practice, and therefore its convergence and generalization properties have been the subject of intense theoretical investigation over the past few years [52, 27, 58, 57, 45, 13, 15, 46]. Many variants of Local SGD exist, including those that use random reshuffling instead of i.i.d. sampling locally [61, 40], adaptive methods such as Adam [48, 55], and modifications to handle data heterogeneity [26, 41], personalization [17], or additionally use gradient compression [16, 50]. Generalized Local SGD, where we use two stepsizes (as in Algorithm 1), is known to be important in managing the trade-off between converging quickly and converging to a mismatched point in heterogeneous distributed optimization [57, 5, 46]. Our focus here is on the homogeneous or i.i.d. data setting; Here, the most related works are [26, 37, 24, 53] and we discuss our work’s relation to theirs in detail in the next section after reviewing some preliminaries.
3 Theory
In this section we conduct the study our main algorithm, Generalized Local SGD (Algorithm 1 with $\mathrm{LocalUpdate}(y,g)=y-\eta g$ and $\mathrm{OuterUpdate}(x,\Delta)=x-\gamma\Delta$). We first review some preliminaries, then present our main results.
3.1 Preliminaries
We are solving the optimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$, where we assume $f$ satisfies the following curvature and regularity condition.
Assumption 3.1.
The function $f$ is differentiable, convex, has $L$-Lipschitz gradients, and has a minimizer $x_{\ast}$.
We suppose that we can access a stochastic first-order oracle that given a point $x$ returns a gradient $g(x)$ that satisfies the following assumption.
Assumption 3.2.
Given a point $x\in\mathbb{R}^{d}$, the stochastic gradients $g(x)\in\mathbb{R}^{d}$ are (a) unbiased in expectation $\mathbb{E}\left[g(x)\right]=\nabla f(x)$, and (b) has variance bounded as $\mathbb{E}\left[\left\lVert g(x)-\nabla f(x)\right\rVert^{2}\right]\leq\sigma^{2}$, where $\mathbb{E}\left[\cdot\right]$ denotes the expectation operator.
Our setting is distributed, but with identically distributed data: there are $M$ different nodes, but they all sample stochastic gradients from the same data distribution in an i.i.d. (independent and identically distributed) manner. We denote the inner product between two vectors $a$ and $b$ by $\left\langle a,b\right\rangle$ and by $\left\lVert\cdot\right\rVert$ the corresponding Euclidean norm. For the purpose of theoretical analysis, can write Generalized Local SGD succinctly as
$\vskip-2.84526pty_{m,r,0}$ $=x_{r},\qquad g_{m,r,h}=\text{Stochastic gradient of $y_{m,r,h}$}$
[…]
$x_{r+1}$ $=x_{r}-\gamma\eta\sum_{h=0}^{H-1}\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}.$ (GEN-LOC-SGD)
To simplify our analysis, we follow [52] and define the virtual sequences
$\vskip-2.84526pt\begin{split}y_{r,h}\overset{\text{def}}{=}\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h},\qquad\qquad g_{r,h}\overset{\text{def}}{=}\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}\\ \overline{g}_{m,r,h}\overset{\text{def}}{=}\mathbb{E}_{r,h-1}\left[g_{m,r,h}\right]=\nabla f(y_{m,r,h}),\qquad\qquad\overline{g}_{r,h}\overset{\text{def}}{=}\mathbb{E}_{r,h-1}\left[g_{r,h}\right].\end{split}$ (1)
3.2 Main convergence result
Recall that we consider Algorithm 1 the particular case when $\mathrm{LocalUpdate}(y,g)=y-\eta g$ and $\mathrm{OuterUpdate}(x,\Delta)=x-\gamma\Delta$.
Existing results on the convergence of Gen. Local SGD. When the outer stepsize $\gamma=1$, the convergence of (GEN-LOC-SGD) is very well understood, with tightly matching upper and lower bounds [27, 58, 13]. In particular, the best rate for the algorithm is
$\mathbb{E}\!\left[\!f\!\left(\!\frac{1}{RH}\!\sum_{r=0}^{R-1}\!\sum_{h=0}^{H-1}\!y_{r,h}\!\right)\!\right]\!-f(x_{\ast})\leq\mathcal{O}\!\left(\!\frac{L{\left\lVert x_{0}\!-x_{\ast}\right\rVert}^{2}}{RH}+\frac{\sigma\left\lVert x_{0}\!-x_{\ast}\right\rVert}{\sqrt{MRH}}+\frac{L^{\frac{1}{3}}\sigma^{\frac{2}{3}}\left\lVert x_{0}\!-x_{\ast}\right\rVert^{\frac{4}{3}}}{H^{\frac{1}{3}}R^{\frac{2}{3}}}\!\right).$ (2)
The first two terms in the above convergence guarantee show that increasing the number of local steps has the same effect as increasing the number of communication rounds $R$, and are identical to the convergence guarantee of doing $RH$ steps of SGD with minibatch size $M$. Local SGD differs from ordinary minibatch SGD in the last term, which shows different scaling between $H$ and $R$, where increasing $R$ helps more than increasing $H$. This is because increasing $H$ incurrs additional client drift that slows down the convergence of the algorithm in the presence of stochastic gradient noise. When the outer stepsize $\gamma$ is allowed to vary, the convergence of the algorithm is less clear. [26] gives the following convergence rate in the absence of data heterogeneity,
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}+\frac{\sigma\left\lVert x_{0}-x_{\ast}\right\rVert}{\sqrt{MR}}\right),$
for specially chosen $\eta$ and $\gamma$ pairs. This rate matches that of Minibatch SGD, but does not recover the convergence rate of vanilla Local SGD given by Equation˜2. [24] also give a guarantee for Generalized Local SGD with a specific outer learning rate that is always at least $1$ and that depends on the heterogeneity of the iterates across the different clients. Since the analysis is conducted in the heterogeneous setting, the local stepsize required to scale with $1/H$. A guarantee that applies to any outer learning rate in the nonconvex, heterogeneous setting given by [53].
The limiting factor in existing analysis is that we are forced to choose the local stepsize $\eta$ to scale as $\frac{1}{LH}$, whereas to obtain Equation˜2 we sometimes need to choose $\eta$ to be much larger, on the order of $\frac{1}{L}$. If we aim to accurately characterize the convergence of (GEN-LOC-SGD), our analysis has to encompass both large and small local stepsizes $\eta$.
New analysis. We now present our main convergence theorem for (GEN-LOC-SGD).
Theorem 3.3.
Suppose that Assumptions˜3.1 and 3.2 hold. Then the iterates generated by Generalized Local SGD run with local stepsize $\eta>0$ and outer stepsize $\gamma>0$ for $R$ communication rounds and with $H$ local steps per round satisfy,
$\!\mathbb{E}\!\left[f\!\left(\!\frac{1}{RH}\!\sum_{r=0}^{R-1}\!\sum_{h=0}^{H-1}\!y_{r,h}\!\right)\!\right]\!-f(x_{\ast})\leq\mathcal{O}\left(\!\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma RH}+\frac{\eta\sigma^{2}(1+(\gamma-1)_{+})}{M}+L\eta^{2}\sigma^{2}H\!\right),$ (3)
where $(a)_{+}=\max(a,0)$. and provided the stepsizes $\eta$ and $\gamma$ jointly satisfy $\eta L(1+(\gamma-1)_{+}H)\leq\frac{1}{4}$.
Implications of Theorem˜3.3. Before giving a proof sketch for Theorem˜3.3, we first discuss its implications. Observe that when $\gamma\leq 1$, we are allowed to choose $\eta$ larger than $\Omega(\frac{1}{LH})$. This is crucial to obtain the rate of Equation˜2. Indeed, when $\gamma=1$, the requirement on $\eta$ reduces to $\eta L\leq\frac{1}{4}$ and we can choose $\eta$ following [58] as
$\eta=\min\left(\frac{1}{4L},\sqrt{\frac{M{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\sigma^{2}RH}},\left[\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{L\sigma^{2}H^{2}R}\right]^{\frac{1}{3}}\right)$
Plugging this choice of $\eta$ yields the convergence guarantee of Equation˜2. Alternatively, when $8\eta L\leq 1$, the stepsize requirement is met if we choose $\eta\gamma LH\leq\frac{1}{8}$ and we immediately get the Minibatch SGD guarantee. In particular, choose $\eta\!=\!\mathcal{O}\!\left(\!\frac{1}{RL}\!\right)$ and $\gamma\!=\!\mathcal{O}\!\left(\!\frac{\gamma_{\ast}}{\eta LH}\!\right)$, the rate then becomes
$f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{8L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma_{\ast}R}+\frac{\sigma^{2}H}{8R^{2}L}+\frac{\gamma_{\ast}\sigma^{2}}{4LMH},$
where $y_{\mathrm{out}}$ denotes the average over all iterations and clients as in Equation˜3. Then for $R$ large enough we can choose $\gamma_{\ast}=\mathcal{O}\left(\sqrt{\frac{LD^{2}\sigma^{2}MH}{R\sigma^{2}}}\right)$ and this gives us the minibatch SGD rate
$f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{LD^{2}}{R}+\frac{\sigma D}{\sqrt{MRH}}.$
This confirms the intuition that at the extremes, manipulating the stepsizes $\gamma$ and $\eta$ allows us to interpolate between minibatch SGD and (vanilla) Local SGD, as observed by [57]. In fact, Theorem˜3.3 allows us to go a step further and get an explicit expression for the optimal inner and outer stepsizes depending on the problem parameters. This is given by the following proposition.
Proposition 3.4.
Let $h(\eta,\gamma)$ be defined as
$h(\eta,\gamma)=\frac{D^{2}}{\eta\gamma RH}+L\sigma^{2}H\eta^{2}+\frac{\eta(1+(\gamma-1)_{+})\sigma^{2}}{M}.$ (4)
Consider the optimization problem:
$\min_{\eta>0,\gamma>0}h(\eta,\gamma)\qquad\text{subject to}\quad$ $\eta L\left(1+(\gamma-1)_{+}H\right)\leq\frac{1}{4}.$ (5)
The solution $(\eta^{*},\gamma^{*})$ is given by comparing the following two candidates.
  1. 1.
    Candidate $(\eta_{A}^{*},\gamma_{A}^{*})$ defined by $\gamma_{A}^{*}=1$ and $\eta_{A}^{*}=\min(\frac{1}{4L},\eta_{A}^{\prime})$ where $\eta_{A}^{\prime}$ is the unique positive root of the cubic equation
    $2LH\sigma^{2}\eta^{3}+\frac{\sigma^{2}}{M}\eta^{2}-\frac{D^{2}}{RH}=0.$
  2. 2.
    Candidate $(\eta_{B}^{*},\gamma_{B}^{*})$ for the regime $\gamma\geq 1$ with $4\eta L<1$, where (a) the constraint is enforced with equality:
    $\gamma_{B}(\eta)=1+\frac{1}{H}\left(\frac{1}{4L\eta}-1\right),$
    and (b) $\eta_{B}^{\ast}$ is the unique positive root of the cubic equation
    $-\frac{4L^{2}D^{2}(H-1)}{R}+2L\sigma^{2}H\eta\left(\eta L(H-1)+1\right)^{2}+\frac{\sigma^{2}(H-1)}{MH}(\eta L(H-1)+1)^{2}=0.$
The optimal solution $(\eta^{*},\gamma^{*})$ is the candidate pair from $\{(\eta_{A}^{*},\gamma_{A}^{*}),(\eta_{B}^{*},\gamma_{B}^{*})\}$ that yields the smaller value of $h(\eta,\gamma)$.
The proof of the above proposition is straightforward and follows by writing the KKT conditions for the optimization problem in Equation˜5. A consequence of Proposition˜3.4 is that in the case of ill-tuning of the inner stepsize $\eta$, a large outer stepsize $\gamma$ can make up for it. For example, if $\sigma\rightarrow 0$ and $\eta LH\ll\mathcal{O}(1)$, we can make up for this by choosing $\gamma$ as $\frac{1}{\eta LH}$. Thus, we can interpret the outer learning rate $\gamma$ as having two dual roles. (a) It allows us to interpolate between minibatch SGD $(\gamma>1)$ and vanilla Local SGD $(\gamma=1)$, giving us the better of the two rates, and (b) it provides us some additional leeway in hyperparameter tuning by making up for ill-tuned inner learning rate $\eta$.
Our theory suggests that in the worst case, choices of $\gamma<1$ are not useful from an optimization perspective. We should either choose $\gamma=1$ or $\gamma>1$. This can be seen even on quadratic objectives, for example if $f(x)=\frac{x^{\top}Qx}{2}$ for some positive definite matrix $Q$, then a straightforward computation gives the expected iterate after $H$ local steps and $R$ communication rounds is
$\mathbb{E}\left[x_{R}\right]=((1-\gamma)I+\gamma(I-\eta Q)^{H})x_{0}.$
From this, it is clear that if $\eta$ is chosen such that $(I-\eta Q)^{H}$ has eigenvalues smaller than $1$, we should choose $\gamma\geq 1$. While if $(I-\eta Q)^{H}$ has any eigenvalues larger than $1$, we should just choose $\gamma=0$ (i.e. just don’t apply the algorithm at all). In other words, $\gamma$ can make up for a learning rate that is too small, but not a learning rate that is too large. This observation does not exclude that $\gamma<1$ can be useful from a generalization perspective, as noted for the case of a single client by [62], in the presence of data heterogeneity, as noted by [4], or in the presence of time-varying noise (see next subsection).
Proof sketch for Theorem˜3.3. We first start by expanding the update for the round iterate $x_{r+1}-x_{\ast}=x_{r+1}-x_{r}+x_{r}-x_{\ast}$ similar to [26] to get,
${\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}={\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-2\gamma\eta\sum_{h=0}^{H-1}\left\langle x_{r}-x_{\ast},g_{r,h}\right\rangle+\gamma^{2}\eta^{2}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}$
$\qquad={\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-2\gamma\eta\sum_{h=0}^{H-1}\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle-2\gamma\eta\sum_{h=0}^{H-1}\left\langle y_{r,h}-x_{\ast},g_{r,h}\right\rangle+\gamma^{2}\eta^{2}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2},\vskip-8.53581pt$
where $g_{r,h}$ is defined as in Equation˜1. [26, 24] control the inner product $\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ by either using smoothness or Young’s inequality; This would force us to bound the stray ${\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}$ and take the local stepsize $\eta$ to be small in order to ensure convergence. Instead, we rely on bounding this quantity directly by viewing it as the regret in the online convex optimization sense with respect to the comparator $x_{r}$. Observe that the virtual sequence of averaged local iterates satisfies $y_{r,h+1}=y_{r,h}-\eta g_{r,h}$, and thus through standard regret analysis we have
$\sum_{h=0}^{H-1}-\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ $=\frac{-{\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$ (6)
The negative terms ${\left\lVert y_{r,H}-x_{r}\right\rVert}^{2}$ in Equation˜6 turn out to be crucial in obtaining an analysis that works for all $\eta$ and not just small $\eta$. With this change and through carefully bounding the variance terms following [27, 58], we obtain the guarantee of Theorem˜3.3. The full proof is provided in Section˜B.2.
Comparison with results on related algorithms. [37] analyze a closely related variant of the algorithm that uses federated random reshuffling [40] as a base. This is a significantly different algorithm that doesn’t allow for an arbitrary number of local steps $H$ and depends on $f$ posessing finite-sum structure. Nevertheless, we can still specialize [38, Theorem 2] approximately to our setting, by using $H$ as the number of data points in an epoch. In our notation, their convergence guarantee reads
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+\eta^{2}H^{2}\sigma^{2}\right),$
under the conditions $\eta H\leq\frac{1}{L}$ and $1\leq\gamma\leq\frac{1}{L\eta H}$. Their theory thus also suggests that $\gamma\geq 1$ can be useful. Optimizing over $\eta$ and $\gamma$ yields the convergence rate
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}\right),$
this rate is the same as gradient descent for $R$ steps (since the finite-sum structure means that per-epoch we approximate one step of gradient descent when $\eta$ is small). A similar rate is derived in [30, 31] if we have access to the proximal operator (i.e. we can do many local steps $H$ on a modified objective). [30] in particular show that an outer learning rate greater than $1$ can be particularly useful for improving the convergence of FedProx [32] in the heterogeneous setting when the smoothness constant varies significantly between different clients.
Analysis with momentum. Our analysis suggests that values of $\gamma>1$ are potentially very useful, but in practice such values are rarely used. One reason this might be the case is because the momentum effectively acts as a stepsize multiplier, i.e. in the presence of momentum parameter $\mu$ the effective outer stepsize becomes $\frac{\gamma}{1-\mu}$. Our next theorem establishes this rigorously.
Theorem 3.5.
Suppose that Assumptions˜3.1 and 3.2 hold. Suppose that the outer update is gradient descent with momentum, $\mathrm{OuterUpdate}(x_{r},-\Delta_{r,H})=x_{r}+\gamma\Delta_{r,H}+\mu(x_{r}-x_{r-1})$ with momentum parameter $\mu\in[0,1)$ and the local update is gradient descent $\mathrm{LocalUpdate}(y,g)=y-\eta g$ in Algorithm˜1. Let the step sizes $\eta,\gamma$ satisfy $\eta L\left(1+\left(\frac{\gamma}{1-\mu}-1\right)_{+}H\right)\leq\frac{1}{4}$ and $\frac{\eta\gamma\mu LH}{1-\mu}\leq\frac{1}{16}$. Then after $R$ rounds of communication, the averaged iterate satisfies
$\begin{split}\mathbb{E}\left[f(\overline{y})\right]-f(x_{*})&\leq\mathcal{O}\left(\frac{(1-\mu){\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+L\eta^{2}\sigma^{2}H+\frac{\eta\sigma^{2}\left(1+(\frac{\gamma}{1-\mu}-1)_{+}\right)}{M}+\frac{\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}}{M}\right),\end{split}$
where $\bar{y}$ is defined as the average of all local iterates across training (as in Equation˜3) and $(a)_{+}=\max(a,0)$.
The proof is provided in Section˜B.3. Theorem 3.5 shows the requirement on the outer stepsize is relaxed from a requirement on $\gamma$ to a requirement on $\frac{\gamma}{1-\mu}$, allowing us to reap the same benefits of $\gamma>1$ observed earlier if we also tune $\mu$. This benefit was first observed in [53] for nonconvex optimization with small local stepsize $\eta$ provided we use an additional momentum buffer. Our work gives direct theoretical support to this observation even with a single momentum buffer and allowing for large $\eta$.
3.3 Convergence with accelerated outer optimizer
We now consider the use of acceleration. To the best of our knowledge, the combination of an accelerated outer optimizer with an unaccelerated inner optimizer, as in e.g. DiLoCo [10, 23], has not been analyzed in the literature before. We take steps towards addressing this gap and understanding the convergence properties of such algorithms by considering Nesterov’s accelerated gradient descent [43] as the outer optimizer and (stochastic) gradient descent as the inner optimizer. The following theorem gives a convergence guarantee for this setting.
Theorem 3.6.
Suppose that Assumptions˜3.1 and 3.2 hold and the stepsizes satisfy $2L\eta\leq 1$ and $\gamma\leq 1$. Suppose that the outer update is accelerated gradient descent with Nesterov momentum as follows
$u_{r+1}=x_{r}-\eta\Delta_{r,H},$ $z_{r+1}=z_{r}-\gamma_{r}\eta\Delta_{r,H},$ $x_{r+1}=(1-\tau_{r+1})u_{r+1}+\tau_{r+1}z_{r+1},$
with parameters $\gamma_{r}=\frac{\gamma(r+1)}{2}$ and $\tau_{r}=\frac{2}{r+2}$, and the local update is gradient descent $\mathrm{LocalUpdate}(y,g)=y-\eta g$ in Algorithm˜1. Then after $R$ rounds of $H$ steps, the final iterate $u_{R}$ satisfies
$\mathbb{E}\left[f(u_{R})\right]-f(x_{*})\leq\frac{2{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma\eta R^{2}H}+\frac{RL\eta^{2}\sigma^{2}H}{2M}+\frac{RL^{2}\eta^{3}\sigma^{2}H^{2}}{2}+\frac{\gamma\eta\sigma^{2}R}{2M}.$ (7)
To understand the implications of the above guarantee, we specialize it with a tuned pair of learning rates $(\gamma,\eta)$ below.
Corollary 3.7.
In the same setting as Theorem˜3.6, setting $\gamma=1$ in Equation˜7, and choosing
$\eta=\min\left\{\frac{1}{2L},\left(\frac{2MD^{2}}{R^{3}L\sigma^{2}H^{2}}\right)^{1/3},\left(\frac{4D^{2}}{3R^{3}L^{2}\sigma^{2}H^{3}}\right)^{1/4},\sqrt{\frac{4MD^{2}}{R^{3}H\sigma^{2}}}\right\},$
where $D=\left\lVert x_{0}-x_{\ast}\right\rVert$ and the final iterate $u_{R}$ satisfies
$\mathbb{E}[f(u_{R})]-f(x_{*})$ $\leq\mathcal{O}\Bigl{(}\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RM^{1/3}H^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{R^{5/4}H^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\Bigr{)}.$ (8)
Equation˜8 shows that in the absence of noise, we obtain a rate accelerated in $R$ but not $H$. This intuitively makes sense, since we do acceleration only in the outer loop. In the presence of noise, we have in the worst-case the unimprovable $\frac{\sigma D}{\sqrt{MRH}}$ term and two additional noise terms that characterize the drift suffered by this algorithm. Notably, the drift terms have much better dependence on $R$ compared to Local SGD, as given by Equation˜2. [60] analyze FedAC, an accelerated variant of Local SGD that uses acceleration locally and applies simple averaging as the outer optimizer. Their algorithm enjoys the convergence rate
$\mathbb{E}[f(x_{\mathrm{out}})]-f(x_{\ast})\leq\mathcal{O}\left(\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RH^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{RH^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\right).$
Comparing with Equation˜8, our algorithm enjoys better dependence on $R$ and $M$ in the denominators of the two drift terms while using momentum sequences only on the server.
3.4 Data-dependent convergence result
To further understand the role of the outer stepsize, we now present a data-dependent, high-probability guarantee for Generalized Local SGD in Theorem 3.8, compared to the rather worst-case analysis of Theorem 3.3. This analysis may also provide insights into practical tuning of the outer learning rate
Theorem 3.8.
Suppose that Assumptions˜3.1 and 3.2 hold. Then in Algorithm 1 with outer update $x=x-\gamma\Delta$ and local update $y=y-\eta g$, if the local stepsize satisfies $\eta\leq\frac{1}{L}$ then with probability at least $1-\delta$ the iterates generated satisfy
$\begin{split}&f\left(\frac{1}{RH}\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}y_{r,h}\right)-f(x_{\ast})\leq\tilde{\mathcal{O}}\Biggl{(}\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma\eta RH}+\frac{\gamma\eta}{RH}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}+\gamma\eta\sigma^{2}\\ &+\frac{\left\lvert 1-\gamma\right\rvert\eta}{RH}\sum_{r}\left(\sum_{h}\left\lVert g_{r,h}\right\rVert\right)^{2}+\frac{\eta}{\gamma H}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}+\eta\sigma\sqrt{\frac{1}{MR}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}}\Biggr{)}.\end{split}$
The proof of Theorem 3.8 is provided in Section˜B.5. Compared to Theorem 3.3, the guarantee we obtain here is weaker in some areas, e.g., the variance term $\gamma\eta\sigma^{2}$ does not benefit from increasing $M$. On the other hand, this guarantee is a high-probability and data-dependent guarantee. To the best of our knowledge, this is the first high-probability convergence guarantee for Local SGD in the literature. Theorem 3.8 allows us to observe another potential benefit of using $\gamma\neq 1$. To see how, let us make the simplifying assumption that $\left\lVert\hat{g}_{r,h}\right\rVert\approxeq G_{1}$ and $\left\lVert g_{m,r,h}\right\rVert\approxeq G_{2}$. Observe that by the triangle inequality we have $G_{1}\leq G_{2}$, but in fact $G_{1}$ can be significantly smaller than $G_{2}$, particularly in the later stages of the optimization process, due to the variance reduction effect of averaging together the gradients on different nodes. Then we can rewrite the above guarantee as
$\hskip-2.84526pt\!f\!\left(\overline{y}\right)\!-\!f(x_{\ast})$ $\leq\tilde{\mathcal{O}}\!\left(\!\frac{d_{0}^{2}}{\gamma\eta RH}+\!\gamma\eta G_{1}^{2}+\!\gamma\eta\sigma^{2}+\!\left\lvert 1\!-\!\gamma\right\rvert\eta HG_{1}^{2}+\!\frac{\eta HG_{2}^{2}}{\gamma}+\!\eta\sigma\sqrt{H}G_{2}\!\right)$ (9)
The $\gamma$ that minimizes this upper bound is given by the following proposition.
Proposition 3.9.
Let $g(x)=\frac{a}{x}+bx+\left\lvert 1-x\right\rvert c$ for $a,b,c\geq 0$.
  • if $a\geq b+c$, then $\sqrt{a/(b+c)}$ minimizes $g$,
  • if $b-c\geq 0$ and $a\leq b-c$, then $\sqrt{a/(b-c)}$ minimizes $g$,
  • Otherwise, $x=1$ minimizes $g$.
Applying this lemma to Equation˜9 one can see that simple averaging is suboptimal depending on the variance and relative magnitudes of $G_{1}$ and $G_{2}$. In particular, the first condition in our setting is
$\frac{d_{0}^{2}}{\eta RH}+\eta HG_{2}^{2}\gtrsim\eta(G_{1}^{2}+\sigma^{2})+\eta HG_{1}^{2},$
where $\gtrsim$ indicates that the inequality holds up to constant factors of the terms on both sides. Since $G_{2}\geq G_{1}$, we can simplify the above condition to $\frac{d_{0}^{2}}{\eta^{2}RH}+HG_{2}^{2}\gtrsim\sigma^{2}$. This condition essentially asks if the noise is large relative to the “optimization term” $\frac{d_{0}^{2}}{\eta^{2}RH}$ or not. In the latter case, choosing $\gamma>1$ is helpful, and the outer optimizer acts as a form of momentum that helps reduce the optimization term further. On the other hand, the second condition yields $\gamma<1$ and requires that $\sigma^{2}\gtrsim\frac{d_{0}^{2}}{\eta^{2}RH}+HG_{2}^{2}$. This is an especially noise-dominated regime, which we may expect to observe towards the end of the training process. In this case, decaying the outer learning rate to $\gamma\ll 1$ allows the algorithm to maintain convergence despite the high noise magnitude. When the optimization term and the noise term are of the same order, then $\gamma=1$ is the optimal choice.
4 Experiments
We conduct two sets of experiments: (a) solving convex optimization problems to provide the most direct verification of the predictions of our theory, and (b) training transformer based language models. Due to limitations of space, we present only highlights of the results here and most of the details and ablations are provided in the supplementary materials (Appendix˜A).
(a) Optimal $\gamma$ vs $\sigma$.
(b) Loss for different $\sigma$.
Figure 1: Effect of varying $\sigma$ and $\gamma$ for quadratic problem
4.1 Convex optimization
We conduct experiments on the quadratic objective $f(x)=\frac{1}{2}{\left\lVert Q(x-x_{\ast})\right\rVert}^{2}$, where $Q=A^{\top}A\in\mathbb{R}^{d}$ for $d=50$ and the entries $A_{i,j}$ are all drawn from a normal distribution $A_{i,j}\sim\mathcal{N}(0,1)$ for $i=1,\ldots,d$ and $j=1,\ldots,d$, and $x_{\ast}$ is similarly drawn from the standard $d$-dimensional Gaussian. We use stochastic gradients of the form $g(x)=\nabla f(x)+v$, where the $v$’s are random vectors drawn from the Gaussian with mean 0 and variance $\sigma^{2}$, $v\sim\mathcal{N}(0,\sigma^{2})$. We evaluate the performance of Algorithm 1 for various values of $\sigma$, $\sigma\in\{10^{-3},10^{-2},10^{-1},0.5,1,5,10,15,25,50\}$. For each $\sigma$ we perform an extensive grid search over $\gamma\in\{0.001,0.01,0.1,0.5,0.9,1.0,1.1,1.25,1.5,2\}$ to determine the best one in terms of minimum average loss over the last ten rounds. We use $R=1000$ rounds and $H=50$ local steps, and fix $\eta=0.001$ in all cases.
Figure 1(a) shows how the optimal value of $\gamma$ varies with different noise levels $\sigma$. We observe that, as $\sigma$ increases, the optimal $\gamma$ decreases from $1.0$ to $0.1$, as predicted by our analysis. Figure 1(b) also illustrates the loss trajectories for different noise levels $\sigma$ with the best $\gamma$.
4.2 Transformer pretraining
Setup Following the DiLoCo paper [10], we experiment using a Chinchilla decoder transformer [18] on the C4 dataset [47]. The architecture hyperparameters are identical from the DiLoCo paper [10] and are given in Section˜A.1.1. We fix the batch size at $512$ and the sequence length at $1024$. We experiment at different scales, from 150 million to 1 billion parameters. For all experiments, the inner optimizer is AdamW [36] trained with a cosine learning rate schedule defined across the total amount of steps. The inner optimizer state is never shared across replicas, and is passed from one round to the other.
(a) 150M
(b) 400M
(c) 1B
Figure 2: Scaling distributed pretraining, at 150M, 400M, and 1B parameters.
Methods We compare three distributed methods, using different outer optimizers: SGD(lr=1) (equivalent to simple averaging of local models [39]), Nesterov (equivalent to DiLoCo [10]), and ScheduleFree-SGD (SF-SGD) [8]. We use SF-SGD to substitute for outer learning rate scheduling, though it still requires tuning hyperparameters. We also include two “high-communication" data-parallel baselines: one with the global batch size as the local per-replica batch size used by the distributed methods, and one with the same batch size as the global batch size ($M\times$ the local per-replica batch size) used by the distributed methods. The latter requires either more GPUs and more thus communication, or gradient accumulation and thus more time. The latter also has an equal flops budget as the distributed methods. We tuned all our optimizers on the pretraining setting on a separate validation set . We also considered using SF-Nesterov, but it was hard to tune and unstable.
Hyperparameter Selected Range considered
Number of inner steps H 50, 500 50 to 2000
Peak outer LR for Nesterov 0.7 0.1 to 2.0
Peak outer LR for SF-SGD 2.0 $1e^{-4}$ to 10.0
b1 for SF-SGD 0.2 0.0 to 0.99
Peak inner learning rate (150M) $4e^{-4}$ $4e^{-4}$
Peak inner learning rate (400M) $4e^{-4}$ $4e^{-4}$
Peak inner learning rate (1B) $2e^{-4}$ $2e^{-4}$
Table 1: Optimizer hyperparameters for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [18].
Results Table˜1 gives the optimal hyperparameters per scale, and Figure˜2 gives the perplexity curves. Consistent with the predictions of our theory, we found that an outer learning rate greater than $1.0$ performed best for SF-SGD and a relatively large effective outer learning rate also performed best for Nesterov; Moreover, acceleration consistently improved performance relative to the baseline Local SGD. In the supplementary material, we report the effect of varying the number of local steps (Section˜A.1.2), the number of clients/replicas and different ways of FLOPs allocation (Section˜A.1.3), and gradient variance (Section˜A.1.6).
5 Conclusion and Future Work
In this paper, we studied the impact of the outer learning rate on the convergence of Local SGD through two novel convergence theorems that characterize its role in balancing a trade-off between convergence speed and stochastic gradient variance. We have also studied the impact of using momentum in the presence of an outer learning rate, and provided a new convergence analysis for using Nesterov acceleration in the outer optimizer. One limitation of our results is that we only consider the i.i.d. setting; Studying the impact of data heterogeneity is therefore a natural next step. Another avenue for future work is to investigate the role of adaptive outer optimizers in enhancing robustness to client failures and communication delays.
Supplementary material
Appendix A Supplementary experimental details
In this section we provide the details on the language model pretraining experiments discussed in the main text.
A.1 Language model pretraining
We study the impact of using various outer optimizers on large language model pretraining. We utilized Chinchilla-style decoder transformer architectures [18] trained on the C4 dataset [47], consistent with common practices in large-scale model training [10]. The following subsections detail the specific hyperparameters, variations in training configurations (such as the number of inner steps and replicas/clients), and analyses of optimizer behavior, including learning rate scheduling and observed gradient cosine similarities.
A.1.1 Hyperparameters details
We show in Table 1 the hyperparameters considered and kept, and in Table 2 the architectural hyperparameters. We tuned all our optimizers on a separate validation set. We also considered using the Schedule-Free Optimizer with Nesterov acceleration on top but it was hard to tune and unstable.
Table 2: Model Configuration for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [18].
Hyperparameter 150M 400M 1B
Number of layers 12 12 24
Hidden dim 896 1536 2048
Number of heads 16 12 16
K/V size 64 128 128
Vocab size 32,00032{,}000
A.1.2 Varying inner steps
In Figure 3, we compare the stability of different outer optimizers when varying the synchronization frequency. We experiments a different amount of inner steps, from 50, to 2000. All experiments are run in pretraining from scratch, with 150 millions (150M) parameters. We note that as the synchronization frequency decreases (number of inner/local steps increases), performance decreases. Notably, averaging (in orange), is relatively constant w.r.t the synchronization frequency: its performance stay stable from $H=250$ to $H=2000$. On the other hand, using Nesterov with high outer learning rate (in light green) is particularly unstable, its performance decreases by $10.7\%$, this indicates that the learning rate should be tuned alongside the synchronization frequency. On the hand, SF-SGD (in blue) has minimal degradation of performance ($4.2\%$), highlighting the schedule-free property when varying hyperparameters.
Figure 3: Varying the communication frequency, i.e. number of inner steps $H$, when pretraining from scratch at 150M parameters.
Figure 4: Pareto front of the flops vs perplexity, comparing various approach scaling the flops budget: increasing the number of steps, increasing the batch size in data-parallel, and increasing the number of replicas for federated learning.
A.1.3 Varying replicas / flops budget
When increasing the number of distributed replicas, two options are possible: (a) Keeping the local per-replica batch size constant and thus increasing global batch size and flops budget, and (b) Keeping the global batch size/flops budget constant and thus reducing the local per-replica batch size.
We present in Figure 4 results of the first option with x-axis the flops budget for a single model size (150M). It is worth noting that increasing the number of replicas improves the performance of Nesterov (in green) and SF-SGD (in blue) but the gain quickly plateau. On the other hand, increasing the batch size for data-parallel (at the cost of more communication, because more DP replicas) or the number of steps (at the cost of longer training) still rapidly improves perplexity. Therefore, we wish to highlight here a disadvantage of federated learning methods seldom mentioned: while those methods are extremely communication-efficient, and can be made flops-efficient, their flops-efficiency disappear as the number of replicas increases.
To this problem, several hypotheses could be raised, such as the decreasing cosine similarity between outer gradients as the number of replicas increase, even when using an i.i.d. data split across replicas. In Figure 5, we report the average similarity across a whole training for different number of replicas. For momentum-based methods (Nesterov, SF-SGD), the similarity decreases from 30% at $M=2$ replicas to 10% at $M=16$ replicas. Full details across training steps can be found in the appendix.
Finally, note that we didn’t investigate further the second option of keeping the global batch size/flops budget constant and thus reducing the local per-replica batch size. We found that dividing the batch size by the number of replicas leads quickly to a local per-replica batch size that is critically low, and further reduces the flops-efficiency. More investigations should be pushed in that direction.
Figure 5: Cosine similarity between outer gradients across different number of replicas (left) and model scales (right). We average the similarity across the middle 50% of the training.
A.1.4 Schedule-free but not tuning-free
The schedule-free method of [8] enables not doing any learning rate scheduling, greatly simplifying training configuration. However, it doesn’t mean it is hyperparameters-tuning-free. Indeed, we found out that we had to extensively tune the initial learning rate (to $2.0$), remove learning rate warm-up contrarily to what is advised, and use a particularly low $b1$ decay: $0.2$, as illustrated in Figure 7.
Figure 6: Tuning b1 decay has a major impact on performance, and its value must be very low.
Figure 7: Which outer learning rate schedule to use?
A.1.5 Pretraining: outer learning rate scheduling
Schedule-free SGD enables not having to manually scheduling the outer learning rate. Therefore, we wondered if we could improve the SotA federated learning baseline, DiLoCo (Nesterov outer optimizer), with an outer learning rate schedule. We investigate in Figure 7 three schedules: constant as in [10], cosine decay, and linear after a plateau. For the latter we consider a constant plateau for 10% and 25% of the total steps. For each method, we also tuned the peak outer learning rate. We don’t use any warm-up in the outer optimization as we always found it to be harmful.
We find that constant outer learning rate is the best performing schedule. It’s unclear how the other schedules are interacting with the inner learning rate scheduling. A possible solution, not investigated in this report, would be to increase the number of inner steps $H$ as the inner learning rate decreases [14].
A.1.6 Cosine similarity between outer gradients
We display the cosine similarity between outer gradients, across scales (150M, 400M, and 1B) in Figure 8, and across replicas (for 150M, from 2 to 16 replicas) in Figure 9. The solid line represent the mean, and the shaded area the standard deviation. We normalize the x-axis as a percentage of the training in order to compare models which have done different amount of steps (e.g. 24,00024{,}000 steps for 150M vs 30,00030{,}000 for 400M).
(a) 150M
(b) 400M
(c) 1B
Figure 8: Similarity between outer gradients across steps and scales.
(a) M=2 replicas
(b) M=4 replicas
(c) M=8 replicas
(d) M=16 replicas
Figure 9: Cosine similarity between outer gradients across steps and number of replicas.
Theory
Appendix B Guarantees for Local SGD
First, we recall our setting and define some notation. We consider the problem of minimizing a function $f$ in a distributed setting with $M$ workers performing Local SGD. Let $x_{r}$ denote the global model parameters at the beginning of round $r$. Each worker $m$ initializes its local parameters as $y_{m,r,0}=x_{r}$ and performs $H$ local SGD steps according to
$y_{m,r,h+1}=y_{m,r,h}-\eta g_{m,r,h},$
where $g_{m,r,h}=\nabla f(y_{m,r,h})+n_{m,r,h}$ is the stochastic gradient with noise $n_{m,r,h}$, and $\overline{g}_{m,r,h}=\nabla f(y_{m,r,h})$ is the true gradient. By Assumption 3.2 we have $\mathbb{E}\left[g_{m,r,h}\right]=\overline{g}_{m,r,h}$. After $H$ local steps, the global model update can be equivalently written as $x_{r+1}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ where $g_{r,h}=\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}$ is the average gradient across workers and $y_{r,h}=\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h}$ is the average model. Note that these two last sequences are virtual sequences and not actually computed. We also define $x_{r,h}=x_{r}-\gamma\eta\sum_{h=0}^{H-1}g_{r,h}$ as an intermediate quantity used in the analysis.
B.1 Algorithm-independent results
Lemma B.1.
[26, Lemma 6] Let $f$ be a convex and $L$-smooth function. Suppose that $\eta\leq\frac{2}{L}$, let $T_{\eta}(x)=x-\eta\nabla f(x)$. Then
${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}.$
Proof.
The proof is provided for completeness only. We have
${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}$ $={\left\lVert x-y\right\rVert}^{2}+\eta^{2}{\left\lVert\nabla f(x)-\nabla f(y)\right\rVert}^{2}-2\eta\left\langle x-y,\nabla f(x)-\nabla f(y)\right\rangle.$ (10)
By the Baillon-Haddad theorem [1] we have
$\left\langle x-y,\nabla f(x)-\nabla f(y)\right\rangle\geq\frac{1}{L}{\left\lVert\nabla f(x)-\nabla f(y)\right\rVert}^{2}.$
Using this in Equation˜10 gives
${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}-\eta\left(\frac{2}{L}-\eta\right){\left\lVert\nabla f(x)-\nabla f(y)\right\rVert}^{2}.$
If $\eta\leq\frac{1}{L}$ then $\frac{2}{L}-\eta\geq 0$ and therefore ${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}$. ∎
Lemma B.2.
Let $y_{1},\ldots,y_{n}$ be real numbers. Then,
$\frac{1}{n}\sum_{k=1}^{n}\left\lvert y_{i}\right\rvert\leq\sqrt{\frac{1}{n}\sum_{k=1}^{n}y_{i}^{2}}.$
Proof.
This is just the arithmetic mean-root mean square inequality and we include the proof solely for completeness. Let $Y$ be a random variable that takes the value $y_{i}^{2}$ with probability $\frac{1}{n}$, and let $g(x)=\sqrt{x}$. Observe that
$\frac{1}{n}\sum_{k=1}^{n}\left\lvert y_{i}\right\rvert=\mathbb{E}\left[g(Y)\right].$
Since $g$ is a concave function, by Jensen’s inequality we have that $\mathbb{E}\left[g(Y)\right]\leq g(\mathbb{E}\left[Y\right])$. Therefore,
$\frac{1}{n}\sum_{k=1}^{n}\left\lvert y_{i}\right\rvert=\mathbb{E}\left[g(Y)\right]\leq g(\mathbb{E}\left[Y\right])=\sqrt{\frac{1}{n}\sum_{k=1}^{n}y_{i}^{2}}.$
Lemma B.3.
(Variance of Sum of Conditionally Independent Random Variables). Let $Z_{1},\ldots,Z_{n}$ be random variables such that $Z_{i}$ satisfies
$\mathbb{E}_{i-1}\left[Z_{i}\right]=0,\qquad\qquad\text{ and, }$ $\mathbb{E}\left[\left\lVert Z_{i}\right\rVert^{2}\right]=\sigma_{i}^{2},$
where $\mathbb{E}_{i}\left[\cdot\right]$ denotes expectation conditional on $Z_{1},Z_{2},\ldots,Z_{i}$. Then,
$\mathbb{E}\left[\left\lVert\sum_{i=1}^{n}Z_{i}\right\rVert^{2}\right]=\sum_{i=1}^{n}\sigma_{i}^{2}.$
Proof.
$\mathbb{E}\left[\left\lVert\sum_{i=1}^{n}Z_{i}\right\rVert^{2}\right]$ $=\mathbb{E}\left[\mathbb{E}_{n-1}\left[\left\lVert\sum_{i=1}^{n}Z_{i}\right\rVert^{2}\right]\right]$
$=\mathbb{E}\left[\mathbb{E}_{n-1}\left[{\left\lVert\sum_{i=1}^{n-1}Z_{i}\right\rVert}^{2}+{\left\lVert Z_{n}\right\rVert}^{2}+2\left\langle\sum_{i=1}^{n-1}Z_{i},Z_{n}\right\rangle\right]\right]$
$=\mathbb{E}\left[\mathbb{E}_{n-1}\left[\left\lVert\sum_{i=1}^{n-1}Z_{i}\right\rVert^{2}\right]+\sigma_{n}^{2}\right].$
The cross-term $\mathbb{E}_{n-1}\left[2\left\langle\sum_{i=1}^{n-1}Z_{i},Z_{n}\right\rangle\right]$ vanishes because $\mathbb{E}_{n-1}\left[Z_{n}\right]=0$ and $\sum_{i=1}^{n-1}Z_{i}$ is measurable with respect to the sigma-algebra generated by $Z_{1},\ldots,Z_{n-1}$. Continuing,
$\mathbb{E}\left[\left\lVert\sum_{i=1}^{n}Z_{i}\right\rVert^{2}\right]$ $=\mathbb{E}\left[\left\lVert\sum_{i=1}^{n-1}Z_{i}\right\rVert^{2}\right]+\sigma_{n}^{2}.$
Recursing we get,
$\mathbb{E}\left[\left\lVert\sum_{i=1}^{n}Z_{i}\right\rVert^{2}\right]$ $=\sum_{i=1}^{n}\sigma_{i}^{2}.$
This completes the proof. ∎
Lemma B.4.
[21, Lemma 7]. Let $S$ be the set of nonnegative and nondecreasing sequences. Let $y_{1},y_{2},\ldots$ be a sequence in $S$. Let $C_{t}\in\mathcal{F}_{t-1}$ for all $t=1,2,\ldots,T$ and let $X_{t}$ be a martingale difference sequence adapted to $\mathcal{F}_{t}$ such that $\left\lvert X_{t}\right\rvert\leq C_{t}$ with probability $1$ for $t=1,2,\ldots,T$. Then for all $\delta\in(0,1)$ and $\hat{X}_{t}\in\mathcal{F}_{t-1}$ such that $\left\lvert\hat{X}_{t}\right\rvert\leq C_{t}$ with probability $1$, we have that with probability at least $1-\delta-\mathrm{Prob}\left(\exists t\leq T\mid C_{t}>c\right)$ that for all $c>0$
$\left\lvert\sum_{i=1}^{t}y_{i}X_{i}\right\rvert\leq 8y_{t}\sqrt{\theta_{t,\delta}\sum_{i=1}^{t}(X_{i}-\hat{X}_{i})^{2}+c^{2}\theta_{t,\delta}^{2}},$
where $\theta_{t,\delta}=\log\frac{60\log 6t}{\delta}$.
Lemma B.5.
Suppose we have
$r_{k+1}\leq(1+a)r_{k}-b\delta_{k}+c$
Then,
$\min_{j}\delta_{j}$ $\leq\frac{r_{0}e^{aK}}{bK}+\frac{c}{b}.$
Proof.
Let $w_{k+1}=\frac{w_{k}}{1+a}$. We have
$w_{k+1}r_{k+1}$ $\leq(1+a)w_{k+1}r_{k}-bw_{k+1}\delta_{k}+cw_{k+1}$
$=w_{k}r_{k}-bw_{k+1}\delta_{k}+cw_{k+1}.$
Telescoping,
$w_{K}r_{K}$ $\leq w_{0}r_{0}-b\sum_{j=0}^{K-1}w_{j+1}\delta_{j}+c\sum_{j=0}^{K-1}w_{j+1}.$
Rearranging,
$\frac{1}{\sum_{j=0}^{K-1}w_{j+1}}\sum_{j=0}^{K-1}w_{j+1}\delta_{j}\leq\frac{w_{0}r_{0}}{b\sum_{j=0}^{K-1}w_{j+1}}+\frac{c}{b}.$
We have $w_{k}=\frac{w_{k-1}}{1+a}=\frac{w_{0}}{(1+a)^{k}}$. Therefore,
$\sum_{j=0}^{K-1}w_{j+1}$ $=\sum_{j=0}^{K-1}\frac{w_{0}}{(1+a)^{k+1}}$
$\geq\sum_{j=0}^{K-1}\frac{w_{0}}{(1+a)^{K}}$
$=\frac{w_{0}K}{(1+a)^{K}}.$
Therefore,
$\frac{1}{\sum_{j=0}^{K-1}w_{j+1}}\sum_{j=0}^{K-1}w_{j+1}\delta_{j}$ $\leq\frac{r_{0}(1+a)^{K}}{bK}+\frac{c}{b}.$
Finally, it remains to use that $1+a\leq e^{a}$. ∎
B.2 Non-adaptive guarantee without momentum
We begin with a lemma that establishes the regret of the local optimizer. Often the regret is measured against the optimal point (like $x_{\ast}$) but here we instead utilize it against the initial point $y_{r,0}=x_{r}$.
Lemma B.6 (Regret against starting point).
For any learning rate $\eta>0$, the inner product between the displacement from the initial average iterate and the average gradient satisfies,
$\sum_{h=0}^{H-1}\langle y_{r,h}-y_{r,0},g_{r,h}\rangle\leq\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}-\frac{1}{2\eta}{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}.$
Proof.
We begin by using that $y_{r,h+1}=y_{r,h}-\eta g_{r,h}$ and expanding the square as
${\left\lVert y_{r,h+1}-y_{r,0}\right\rVert}^{2}$ $={\left\lVert y_{r,h}-\eta g_{r,h}-y_{r,0}\right\rVert}^{2}$
$={\left\lVert y_{r,h}-y_{r,0}\right\rVert}^{2}+\eta^{2}{\left\lVert g_{r,h}\right\rVert}^{2}-2\eta\left\langle y_{r,h}-y_{r,0},g_{r,h}\right\rangle.$
Rearranging to isolate the inner product term, we obtain
$\left\langle y_{r,h}-y_{r,0},g_{r,h}\right\rangle=\frac{{\left\lVert y_{r,h}-y_{r,0}\right\rVert}^{2}-{\left\lVert y_{r,h+1}-y_{r,0}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}{\left\lVert g_{r,h}\right\rVert}^{2}.$
Summing over $h$ from 0 to H1H-1,
$\sum_{h=0}^{H-1}\left\langle y_{r,h}-y_{r,0},g_{r,h}\right\rangle$ $=\sum_{h=0}^{H-1}\left(\frac{{\left\lVert y_{r,h}-y_{r,0}\right\rVert}^{2}-{\left\lVert y_{r,h+1}-y_{r,0}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}{\left\lVert g_{r,h}\right\rVert}^{2}\right)$
$=\frac{1}{2\eta}\sum_{h=0}^{H-1}({\left\lVert y_{r,h}-y_{r,0}\right\rVert}^{2}-{\left\lVert y_{r,h+1}-y_{r,0}\right\rVert}^{2})+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$
The first sum telescopes
$\sum_{h=0}^{H-1}({\left\lVert y_{r,h}-y_{r,0}\right\rVert}^{2}-{\left\lVert y_{r,h+1}-y_{r,0}\right\rVert}^{2})$ $={\left\lVert y_{r,0}-y_{r,0}\right\rVert}^{2}-{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}$
$=-{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}.$
Therefore,
$\sum_{h=0}^{H-1}\left\langle y_{r,h}-y_{r,0},g_{r,h}\right\rangle$ $=-\frac{{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}$
$\leq\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}-\frac{{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}}{2\eta}.$
Lemma B.7.
(Local client drift bound). Suppose that Assumptions˜3.1 and 3.2 hold. Then in Algorithm GEN-LOC-SGD for all $r$ and $h$, if $\eta\leq\frac{1}{L}$, then
$\mathbb{E}\left[\frac{1}{M^{2}}\sum_{m,s=1}^{M}{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}\right]\leq 2\eta^{2}\sigma^{2}h.$
Proof.
Let $\tilde{T}_{\eta}(y_{m,r,h})=y_{m,r,h}-\eta g_{m,r,h}$ where $g_{m,r,h}$ is the stochastic gradient, and $T_{\eta}(y_{m,r,h})=y-\eta\overline{g}_{m,r,h}$ is the corresponding expected gradient update. We have
$y_{m,r,h+1}-y_{s,r,h+1}$ $=\tilde{T}_{\eta}(y_{m,r,h})-\tilde{T}_{\eta}(y_{s,r,h})$
$=T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})+\left[\tilde{T}_{\eta}(y_{m,r,h})-\tilde{T}_{\eta}(y_{s,r,h})-(T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}))\right]$
$=T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})+\left[\xi_{m,r,h}-\xi_{s,r,h}\right],$
where $\xi_{m,r,h}=\tilde{T}_{\eta}(y_{m,r,h})-T_{\eta}(y_{m,r,h})=-\eta n_{m,r,h}$ is the noise term. Define $\mathcal{V}_{r,h}=\frac{1}{M^{2}}\sum_{m,s=1}^{M}{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}$. It follows that
$\mathcal{V}_{r,h+1}$ $=\frac{1}{M^{2}}\sum_{m,s=1}^{M}{\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}$
$\begin{split}&=\frac{1}{M^{2}}\sum_{m,s=1}^{M}\Bigg{[}{\left\lVert T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})\right\rVert}^{2}+{\left\lVert\xi_{m,r,h}-\xi_{s,r,h}\right\rVert}^{2}\\ &\qquad+2\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),\xi_{m,r,h}-\xi_{s,r,h}\right\rangle\Bigg{]}.\end{split}$
Taking conditional expectation gives
$\begin{split}\mathbb{E}_{r,h}\left[\mathcal{V}_{r,h+1}\right]&=\frac{1}{M^{2}}\sum_{m,s=1}^{M}\Bigg{[}{\left\lVert T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})\right\rVert}^{2}+\mathbb{E}_{h}\left[\left\lVert\xi_{m,r,h}-\xi_{s,r,h}\right\rVert^{2}\right]\Bigg{]}.\end{split}$
Finally, using the fact that ${\left\lVert T_{\eta}(x)-T_{\eta}(y)\right\rVert}^{2}\leq{\left\lVert x-y\right\rVert}^{2}$ whenever $\eta\leq\frac{2}{L}$ (Lemma B.1) and Assumption 3.2, we get
$\mathbb{E}_{r,h}\left[\mathcal{V}_{r,h+1}\right]$ $\leq\frac{1}{M^{2}}\sum_{m,s=1}^{M}\left[{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}+2\eta^{2}\sigma^{2}\right]$
$=\mathcal{V}_{r,h}+2\eta^{2}\sigma^{2}.$
Therefore by taking unconditional expectation and recursing from $h=0$ where all local iterates are equal to $x_{r}$ (so $\mathcal{V}_{r,0}=0$), we get $\mathbb{E}\left[\mathcal{V}_{r,h}\right]\leq 2\eta^{2}\sigma^{2}h$. ∎
Proof of Theorem˜3.3.
W begin by analyzing how the squared distance to the optimal solution changes after one round of communication. From the update rule, we have,
${\left\lVert x_{r+1}-x_{*}\right\rVert}^{2}$ $={\left\lVert x_{r}-x_{*}\right\rVert}^{2}-2\eta\gamma\sum_{h=0}^{H-1}\left\langle x_{r}-x_{*},g_{r,h}\right\rangle+\eta^{2}\gamma^{2}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}.$ (11)
We rewrite the inner product term as
$-\left\langle x_{r}-x_{*},g_{r,h}\right\rangle$ $=\left\langle x_{*}-x_{r},g_{r,h}\right\rangle$
$=\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle+\left\langle y_{r,h}-x_{r},g_{r,h}\right\rangle.$
Summing over all local steps we obtain
$-\sum_{h=0}^{H-1}\left\langle x_{r}-x_{*},g_{r,h}\right\rangle$ $=\sum_{h=0}^{H-1}\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle+\sum_{h=0}^{H-1}\left\langle y_{r,h}-x_{r},g_{r,h}\right\rangle.$
Applying Lemma˜B.6 we get
$-\sum_{h=0}^{H-1}\left\langle x_{r}-x_{*},g_{r,h}\right\rangle$ $=\sum_{h=0}^{H-1}\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle-\frac{{\left\lVert y_{r,H}-y_{r,0}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$ (12)
Observe that since $y_{r,H}-y_{r,0}=-\eta\sum_{h=0}^{H-1}g_{r,h}$, Equation˜12 becomes,
$-\sum_{h=0}^{H-1}\left\langle x_{r}-x_{*},g_{r,h}\right\rangle$ $=\sum_{h=0}^{H-1}\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle-\frac{\eta}{2}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$
Plugging this back into Equation˜11,
${\left\lVert x_{r+1}-x_{*}\right\rVert}^{2}$ $\leq{\left\lVert x_{r}-x_{*}\right\rVert}^{2}+2\eta\gamma\sum_{h=0}^{H-1}\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle$
$\quad+\gamma\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+\eta^{2}\gamma(\gamma-1){\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}.$
Let us take expectation conditional on $x_{1},\ldots,x_{r}$,
$\begin{split}\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+2\eta\gamma\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},g_{r,h}\right\rangle\right]\\ &+\gamma\eta^{2}\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\lVert g_{r,h}\right\rVert^{2}\right]+\eta^{2}\gamma(\gamma-1)\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right].\end{split}$ (13)
For the squared norm of the average gradient:
$\mathbb{E}_{r}\left[\left\lVert g_{r,h}\right\rVert^{2}\right]$ $=\mathbb{E}_{r}\left[\mathbb{E}_{r,h-1}\left[\left\lVert g_{r,h}\right\rVert^{2}\right]\right]$
$=\mathbb{E}_{r}\left[\mathbb{E}_{r,h-1}\left[\left\lVert g_{r,h}-\overline{g}_{r,h}\right\rVert^{2}\right]+{\left\lVert\overline{g}_{r,h}\right\rVert}^{2}\right]$
$=\frac{\sigma^{2}}{M}+\mathbb{E}_{r}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right],$
where we use $\mathbb{E}_{r,h-1}\left[\cdot\right]$ to denote expectation conditional on the $\sigma$-algebra generated by all the stochastic gradients up to and including step h1h-1. Substituting this into Equation˜13,
$\begin{split}\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+2\eta\gamma\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},g_{r,h}\right\rangle\right]+\frac{\gamma\eta^{2}H\sigma^{2}}{M}\\ &+\gamma\eta^{2}\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right]+\eta^{2}\gamma(\gamma-1)\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right].\end{split}$ (14)
Now we bound the inner product term:
$\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},g_{r,h}\right\rangle\right]$ $=\mathbb{E}_{r}\left[\mathbb{E}_{h-1}\left[\left\langle x_{\ast}-y_{r,h},g_{r,h}\right\rangle\right]\right]$
$=\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},\overline{g}_{r,h}\right\rangle\right]$
$=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},\overline{g}_{m,r,h}\right\rangle\right]$
$=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{m,r,h}+y_{m,r,h}-y_{r,h},\overline{g}_{m,r,h}\right\rangle\right]$
$=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{m,r,h},\overline{g}_{m,r,h}\right\rangle\right]+\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle y_{m,r,h}-y_{r,h},\overline{g}_{m,r,h}\right\rangle\right].$
Using Young’s inequality for the second term,
$\mathbb{E}_{r}\left[\left\langle x_{*}-y_{r,h},g_{r,h}\right\rangle\right]$ (15)
$\quad\leq\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle x_{*}-y_{m,r,h},\overline{g}_{m,r,h}\right\rangle\right]+\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\frac{{\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}^{2}}{2\alpha}+\frac{\alpha}{2}{\left\lVert\overline{g}_{m,r,h}\right\rVert}^{2}\right]$
$\quad=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\langle x_{*}-y_{m,r,h},\overline{g}_{m,r,h}\right\rangle\right]+\frac{V_{r,h}}{2\alpha}+\frac{\alpha}{2M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\lVert\overline{g}_{m,r,h}\right\rVert^{2}\right],$ (16)
where $V_{r,h}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\lVert y_{m,r,h}-y_{r,h}\right\rVert^{2}\right]$ by definition. By the convexity of $f$,
$\left\langle x_{\ast}-y_{m,r,h},\overline{g}_{m,r,h}\right\rangle$ $=\left\langle x_{\ast}-y_{m,r,h},\nabla f(y_{m,r,h})\right\rangle$
$\leq f(x_{\ast})-f(y_{m,r,h})$
$=-\left(f(y_{m,r,h})-f(x_{\ast})\right).$ (17)
For the variance term, when $\eta\leq\frac{1}{L}$ we use Lemma˜B.7
$V_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\lVert y_{m,r,h}-y_{r,h}\right\rVert^{2}\right]$
$\leq\frac{1}{M}\sum_{m=1}^{M}\frac{1}{M}\sum_{s=1}^{M}\mathbb{E}_{r}\left[\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert^{2}\right]$
$=\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\mathbb{E}_{r}\left[\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert^{2}\right]$
$\leq 2\eta^{2}\sigma^{2}h\leq 2\eta^{2}\sigma^{2}H.$ (18)
By smoothness,
${\left\lVert\overline{g}_{m,r,h}\right\rVert}^{2}={\left\lVert\nabla f(y_{m,r,h})\right\rVert}^{2}\leq 2L(f(y_{m,r,h})-f(x_{*})).$ (19)
Plugging Equations˜17, 18 and 19 back into Equation˜16 we get
$\mathbb{E}_{r}\left[\left\langle x_{\ast}-y_{r,h},g_{r,h}\right\rangle\right]\leq\frac{-(1-\alpha L)}{M}\sum_{m=1}^{M}(\mathbb{E}_{r}\left[f(y_{m,r,h})\right]-f(x_{\ast}))+\frac{\eta^{2}\sigma^{2}H}{\alpha}.$ (20)
Substituting (20) back into our main recursion (Equation˜13),
$\begin{split}\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-\frac{2\eta\gamma(1-\alpha L)}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}(\mathbb{E}_{r}\left[f(y_{m,r,h})\right]-f(x_{\ast}))+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}\\ &+\frac{\gamma\eta^{2}H\sigma^{2}}{M}+\gamma\eta^{2}\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right]+\eta^{2}\gamma(\gamma-1)\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right].\end{split}$ (21)
We now have two cases. Case 1. If $\gamma\geq 1$, then we have by Lemma˜B.3 and Jensen’s inequality applied to ${\left\lVert\cdot\right\rVert}^{2}$,
$\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right]$ $=\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}(g_{r,h}-\mathbb{E}_{r}\left[g_{r,h}\right])\right\rVert^{2}\right]+{\left\lVert\sum_{h=0}^{H-1}(\mathbb{E}_{r}\left[g_{r,h}\right])\right\rVert}^{2}$
$=\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}(g_{r,h}-\mathbb{E}_{r}\left[g_{r,h}\right])\right\rVert^{2}\right]+{\left\lVert\sum_{h=0}^{H-1}(\mathbb{E}_{r}\left[\mathbb{E}_{r,h-1}\left[g_{r,h}\right]\right])\right\rVert}^{2}$
$=\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}(g_{r,h}-\mathbb{E}_{r}\left[g_{r,h}\right])\right\rVert^{2}\right]+{\left\lVert\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\overline{g}_{r,h}\right]\right\rVert}^{2}$
$\leq\frac{\sigma^{2}H}{M}+\mathbb{E}_{r}\left[\left\lVert\sum_{h=0}^{H-1}\overline{g}_{r,h}\right\rVert^{2}\right]$
$\leq\frac{\sigma^{2}H}{M}+H\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right].$ (22)
Using Jensen’s inequality and smoothness we have
$\mathbb{E}_{r}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right]$ $=\mathbb{E}_{r}\left[\left\lVert\frac{1}{M}\sum_{m=1}^{M}\nabla f(y_{m,r,h})\right\rVert^{2}\right]$
$\leq\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[\left\lVert\nabla f(y_{m,r,h})\right\rVert^{2}\right]$
$\leq\frac{2L}{M}\sum_{m=1}^{M}\mathbb{E}_{r}\left[f(y_{m,r,h})-f(x_{\ast})\right].$ (23)
Using Equations˜22 and 23 into Equation˜21 we get
$\begin{split}&\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}\\ &\quad-\frac{2\eta\gamma(1-\alpha L)-2L\gamma\eta^{2}(1+(\gamma-1)H)}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}(\mathbb{E}_{r}\left[f(y_{m,r,h})\right]-f(x_{\ast}))+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}\\ &\qquad+\frac{\gamma^{2}\eta^{2}H\sigma^{2}}{M}.\end{split}$
$\begin{split}&={\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-\frac{2\eta\gamma\left[1-\alpha L-L\eta(1+(\gamma-1)H)\right]}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}(\mathbb{E}_{r}\left[f(y_{m,r,h})\right]-f(x_{\ast}))\\ &\qquad+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}+\frac{\gamma^{2}\eta^{2}H\sigma^{2}}{M}.\end{split}$
$\begin{split}&={\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-2\eta\gamma H(1-\alpha L-L\eta(1+(\gamma-1)H))\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}+\frac{\eta^{2}\gamma^{2}H\sigma^{2}}{M},\end{split}$ (24)
where in the last line we defined
$\hat{\delta}_{r+1}$ $=\frac{1}{MH}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left(f(y_{m,r,h})-f(x_{\ast})\right)$ (25)
Case 2. If $\gamma\leq 1$, then we can simply drop the last term in Equation˜21 and use Equation˜19 to get
$\begin{split}\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-\frac{2\eta\gamma(1-\alpha L-\eta L)}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left(\mathbb{E}_{r}\left[f(y_{m,r,h})\right]-f(x_{\ast})\right)\\ &\qquad+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}+\frac{\gamma\eta^{2}H\sigma^{2}}{M}\end{split}$
$\begin{split}&={\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-2\eta\gamma H(1-\alpha L-\eta L)\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}+\frac{\gamma\eta^{2}H\sigma^{2}}{M},\end{split}$ (26)
where in Equation˜26 we again used the definition in Equation˜25. Looking at both Equations˜24 and 26 and taking the maximum we get that for any $\gamma$,
$\begin{split}\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-2\eta\gamma H(1-\alpha L-\eta L(1+(\gamma-1)_{+}H))\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]\\ &\qquad+\frac{2\eta^{3}\gamma\sigma^{2}H^{2}}{\alpha}+\frac{\eta^{2}\max\{\gamma^{2},\gamma\}H\sigma^{2}}{M},\end{split}$
where $(x)_{+}=\max(x,0)$ is the ReLU function. Putting $\alpha=\frac{1}{2L}$ we get
$\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]$ $\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-\eta\gamma H(1-2\eta L(1+(\gamma-1)_{+}H))\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]$
$\qquad+4L\eta^{3}\gamma\sigma^{2}H^{2}+\frac{\eta^{2}\max\{\gamma^{2},\gamma\}H\sigma^{2}}{M}.$
Under the requirement that the stepsizes $\eta,\gamma$ satisfy
$\eta L(1+(\gamma-1)_{+}H)$ $\leq\frac{1}{4},$
we obtain our recursion
$\mathbb{E}_{r}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]$ $\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}-\frac{\eta\gamma H}{2}\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]+4L\eta^{3}\gamma\sigma^{2}H^{2}+\frac{\eta^{2}\max\{\gamma^{2},\gamma\}H\sigma^{2}}{M}.$
Taking unconditional expectations and rearranging we obtain,
$\mathbb{E}\left[\hat{\delta}_{r+1}\right]\leq\frac{2}{\gamma\eta H}\left[\mathbb{E}\left[\left\lVert x_{r}-x_{\ast}\right\rVert^{2}\right]-\mathbb{E}\left[\left\lVert x_{r+1}-x_{\ast}\right\rVert^{2}\right]\right]+8L\eta^{2}\sigma^{2}H+\frac{2\eta\max(\gamma,1)\sigma^{2}}{M}.$
Summing up both sides as $r$ varies from 0 to R1R-1 and dividing by $1/R$ we get
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]$ $\leq\frac{2}{\gamma\eta RH}\left[{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}-\mathbb{E}\left[\left\lVert x_{R}-x_{\ast}\right\rVert^{2}\right]\right]+8L\eta^{2}\sigma^{2}H+\frac{2\eta\max(\gamma,1)\sigma^{2}}{M}.$
Observe that we can write $\max(\gamma,1)=1+(\gamma-1)_{+}$. Dropping the negative term and using Jensen’s inequality gives
$\mathbb{E}\left[f\left(\frac{1}{MRH}\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}\sum_{m=1}^{M}f(y_{m,r,h})\right)\right]-f(x_{\ast})\leq\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]$
$\qquad\qquad\leq\frac{2{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma\eta RH}+8L\eta^{2}\sigma^{2}H+\frac{2\eta\max(\gamma,1)\sigma^{2}}{M},$
and this is the statement of our theorem. ∎
B.3 Non-adaptive guarantee with momentum
Proof of Theorem 3.5.
We analyze the momentum variant of Local SGD:
$x_{r+1}$ $=x_{r}-\eta\gamma\left(\sum_{h=0}^{H-1}g_{r,h}\right)+\mu(x_{r}-x_{r-1}).$
Define
$z_{r}$ $=x_{r}+\frac{\mu}{1-\mu}(x_{r}-x_{r-1}).$
Then
$z_{r+1}$ $=z_{r}-\frac{\eta\gamma}{1-\mu}\sum_{h=0}^{H-1}g_{r,h}.$
We have
$\begin{split}&{\left\lVert z_{r+1}-x_{\ast}\right\rVert}^{2}={\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}+\frac{\eta^{2}\gamma^{2}}{(1-\mu)^{2}}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}-\frac{2\eta\gamma}{1-\mu}\sum_{h=0}^{H-1}\left\langle z_{r}-x_{\ast},g_{r,h}\right\rangle\\ &={\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}+\frac{\eta^{2}\gamma^{2}}{(1-\mu)^{2}}{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}-\frac{2\eta\gamma}{1-\mu}\sum_{h=0}^{H-1}\left\langle x_{r}-x_{\ast},g_{r,h}\right\rangle\\ &\qquad-\frac{2\eta\gamma\mu}{1-\mu}\sum_{h=0}^{H-1}\left\langle x_{r}-x_{r-1},g_{r,h}\right\rangle.\end{split}$ (27)
Following the same proof as Theorem˜3.3, we can bound (in expectation)
$\begin{split}-&\frac{2\eta\gamma}{1-\mu}\sum_{h=0}^{H-1}\mathbb{E}_{r}\left[\left\langle x_{r}-x_{\ast},g_{r,h}\right\rangle\right]+\frac{\eta^{2}\gamma^{2}}{(1-\mu)^{2}}\mathbb{E}_{r}\left[{\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert}^{2}\right]\leq-\frac{\eta\gamma H}{2(1-\mu)}\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]\\ &\qquad\qquad+4L\eta^{3}\frac{\gamma}{1-\mu}\sigma^{2}H^{2}+\frac{\eta^{2}H\sigma^{2}}{M}\max\left(\left(\frac{\gamma}{1-\mu}\right)^{2},\frac{\gamma}{1-\mu}\right),\end{split}$ (28)
because the local optimization procedure is the same– the same analysis holds line-by-line, only replacing $\gamma$ by $\frac{\gamma}{1-\mu}$, and requiring instead that
$\eta L\left(1+\left(\frac{\gamma}{1-\mu}-1\right)_{+}H\right)\leq\frac{1}{4}.$ (29)
Using Equation˜28 in Equation˜27 (after taking expectation in the latter) we obtain
$\begin{split}&\mathbb{E}_{r}\left[\left\lVert z_{r+1}-x_{\ast}\right\rVert^{2}\right]\leq{\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}-\frac{\eta\gamma H}{2(1-\mu)}\mathbb{E}_{r}\left[\hat{\delta}_{r+1}\right]+4L\eta^{3}\frac{\gamma\sigma^{2}H^{2}}{1-\mu}\\ &\qquad+\frac{\eta^{2}H\sigma^{2}}{M}\max\left(\left(\frac{\gamma}{1-\mu}\right)^{2},\frac{\gamma}{1-\mu}\right)-\frac{2\eta\gamma\mu}{1-\mu}\sum_{h=0}^{H-1}\left\langle x_{r}-x_{r-1},\overline{g}_{r,h}\right\rangle.\end{split}$ (30)
In the following, we use the shorthand $G_{r}\overset{\text{def}}{=}\sum_{h=0}^{H-1}g_{r,h}$. We now proceed to bound $\sum_{h=0}^{H-1}\left\langle x_{r-1}-x_{r},g_{r,h}\right\rangle=\left\langle x_{r-1}-x_{r},G_{r}\right\rangle$ without using the bounded iterates assumption. We note that by definition:
$x_{r}-x_{r-1}=-\eta\gamma G_{r-1}+\mu(x_{r-1}-x_{r-2}).$
Expanding this out recursively, we get the following formula:
$x_{r}-x_{r-1}=-\eta\gamma\sum_{s=0}^{r-1}\mu^{r-1-s}G_{s}.$
For our analysis, we’ll bound the inner product
$\left\langle x_{r-1}-x_{r},G_{r}\right\rangle$ $=\left\langle\eta\gamma\sum_{s=0}^{r-1}\mu^{r-1-s}G_{s},G_{r}\right\rangle$
$=\eta\gamma\sum_{s=0}^{r-1}\mu^{r-1-s}\left\langle G_{s},G_{r}\right\rangle$
We will actually bound the sum of the momentum terms over $r$, i.e. $\sum_{r}\left\langle x_{r-1}-x_{r},G_{r}\right\rangle$. We have
$\sum_{r}\left\langle x_{r-1}-x_{r},G_{r}\right\rangle$ $=\frac{\eta\gamma}{\mu}\sum_{r}\sum_{s<r}\left\langle\mu^{r-s}G_{s},G_{r}\right\rangle$
$=\frac{\eta\gamma}{2\mu}\left[\sum_{r}\sum_{s}\left\langle\mu^{|r-s|}G_{s},G_{r}\right\rangle-\sum_{r}\parallel G_{r}\parallel^{2}\right].$
To bound the first term above, let $A$ be the $R\times R$ matrix whose $(r,s)$th entry equals $\mu^{|r-s|}$, and let $\Gamma=[G_{1}|G_{2}|\ldots|G_{R}]$. Then
$\sum_{r}\sum_{s}\left\langle\mu^{|r-s|}G_{s},G_{r}\right\rangle=\text{Tr}(\Gamma A\Gamma^{\top}).$
We now apply the Gershgorin circle theorem to bound this sum, observe that largest sum of absolute values of entries in a row satisfy
$1+2\sum_{r=1}^{(R-1)/2}\mu^{r}=1+2\mu\frac{1-\mu^{(R-1)/2}}{1-\mu}=\frac{1+\mu-2\mu^{(R+1)/2}}{1-\mu}\leq\frac{1+\mu}{1-\mu}.$
Then, we have
$\text{Tr}(\Gamma A\Gamma^{\top})\leq\frac{1+\mu}{1-\mu}\text{Tr}(\Gamma\Gamma^{\top})=\frac{1+\mu}{1-\mu}\sum_{r}{\left\lVert G_{r}\right\rVert}^{2}.$
Therefore, taking expectations we have
$-\frac{2\eta\gamma\mu}{1-\mu}$ $\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}\mathbb{E}\left[\left\langle x_{r}-x_{r-1},g_{r,h}\right\rangle\right]=\frac{2\eta\gamma\mu}{1-\mu}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\langle x_{r-1}-x_{r},G_{r}\right\rangle\right]$
$\leq\frac{2\eta\gamma\mu}{1-\mu}\frac{\eta\gamma}{1-\mu}\sum_{r=0}^{R-1}\mathbb{E}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right].$ (31)
Using Lemma˜B.3 we have
$\mathbb{E}\left[\left\lVert\sum_{h=0}^{H-1}g_{r,h}\right\rVert^{2}\right]$ $\leq\frac{\sigma^{2}H}{M}+\mathbb{E}\left[\left\lVert\sum_{h=0}^{H-1}\overline{g}_{r,h}\right\rVert^{2}\right]$
$\leq\frac{\sigma^{2}H}{M}+H\sum_{h=0}^{H-1}\mathbb{E}\left[\left\lVert\overline{g}_{r,h}\right\rVert^{2}\right]$
$\leq\frac{\sigma^{2}H}{M}+2LH^{2}\mathbb{E}\left[\hat{\delta}_{r+1}\right],$
where in the last line we used Jensen’s inequality and smoothness. Using this result in Equation˜31 we get
$-\frac{2\eta\gamma\mu}{1-\mu}$ $\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}\mathbb{E}\left[\left\langle x_{r}-x_{r-1},g_{r,h}\right\rangle\right]\leq\frac{\eta\gamma}{2(1-\mu)}\frac{4\eta\gamma\mu}{1-\mu}\left[\frac{\sigma^{2}RH}{M}+2LH^{2}\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]\right].$ (32)
Rearranging and summing up Equation˜30 then using Equation˜32 we have
$\begin{split}&\mathbb{E}\left[\left\lVert z_{R}-x_{\ast}\right\rVert^{2}\right]\leq{\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}-\frac{\eta\gamma H}{2(1-\mu)}\left[1-\frac{8\eta\gamma\mu LH}{1-\mu}\right]\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]\\ &+4L\eta^{3}\frac{\gamma\sigma^{2}H^{2}}{1-\mu}R+\frac{\eta^{2}H\sigma^{2}}{M}\max\left(\left(\frac{\gamma}{1-\mu}\right)^{2},\frac{\gamma}{1-\mu}\right)R+\frac{\eta\gamma H}{1-\mu}\frac{2\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}R}{M}.\end{split}$
Observe that under the condition
$\frac{\eta\gamma\mu LH}{1-\mu}\leq\frac{1}{16}$
the last inequality becomes
$\begin{split}&\mathbb{E}\left[\left\lVert z_{R}-x_{\ast}\right\rVert^{2}\right]\leq{\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}-\frac{\eta\gamma H}{4(1-\mu)}\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]\\ &+4L\eta^{3}\frac{\gamma\sigma^{2}H^{2}}{1-\mu}R+\frac{\eta^{2}H\sigma^{2}}{M}\max\left(\left(\frac{\gamma}{1-\mu}\right)^{2},\frac{\gamma}{1-\mu}\right)R+\frac{\eta\gamma H}{1-\mu}\frac{2\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}R}{M}.\end{split}$
Continuing the proof and rearranging we get
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\hat{\delta}_{r+1}\right]\leq\frac{4(1-\mu){\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+16L\eta^{2}\sigma^{2}H+\frac{4\eta\sigma^{2}}{M}\max\left(\frac{\gamma}{1-\mu},1\right)+\frac{8\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}}{M}.$
It remains to use Jensen’s inequality. ∎
B.4 Acceleration proofs
We recall the algorithm under analysis as
$y_{m,r,0}$ $=x_{r}\text{ for $m=1,\ldots,M$ }$
$y_{m,r,h+1}$ $=y_{m,r,h}-\eta g_{m,r,h}\text{ for $h=0,1,\ldots,H-1$ }$
$u_{r+1}$ $=x_{r}-\eta\sum_{h=0}^{H-1}g_{r,h}$
$z_{r+1}$ $=z_{r}-\gamma_{r}\eta\sum_{h=0}^{H-1}g_{r,h}$
$x_{r+1}$ $=(1-\tau_{r+1})u_{r+1}+\tau_{r+1}z_{r+1},$
where $g_{r,h}=\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h}$, $\gamma_{r}=\frac{\gamma(r+1)}{2}$, and $\tau_{r}=\frac{2}{r+2}$. Note that under the above, $u_{m,r,h}=y_{m,r,h}$ and $u_{r,h}=\frac{1}{M}\sum_{m=1}^{M}u_{m,r,h}$. We first derive two intermediate lemmas, then proceed to the main proof.
Lemma B.8.
Suppose that the local stepsize $\eta$ satisfies $\eta\leq\frac{1}{2L}$. Then, for all $h\in[H-1]$ and $r$, we have
$Hf(u_{r+1})\leq\frac{1}{M}\sum_{m,h<H}\left[\mathbb{E}\left[f(y_{m,r,h})\right]-\frac{\eta}{4}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}\left[\left\lVert\bar{g}_{m,r,h}\right\rVert^{2}\right]\right]+\frac{L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{\eta^{3}\sigma^{2}H^{3}}{2}.$
Proof.
By smoothness,
$f(u_{r,h+1})$ $\leq f(u_{r,h})+\left\langle\nabla f(u_{r,h}),u_{r,h+1}-u_{r,h}\right\rangle+\frac{L}{2}{\left\lVert u_{r,h+1}-u_{r,h}\right\rVert}^{2}$
$=f(u_{r,h})-\eta\left\langle\nabla f(u_{r,h}),g_{r,h}\right\rangle+\frac{L\eta^{2}}{2}{\left\lVert g_{r,h}\right\rVert}^{2}.$
Taking conditional expectation we have
$\mathbb{E}_{h}\left[f(u_{r,h+1})\right]\leq f(u_{r,h})-\frac{\eta}{M}\sum_{m=1}^{M}\left\langle\nabla f(u_{r,h}),\nabla f(u_{m,r,h})\right\rangle+\frac{L\eta^{2}\sigma^{2}}{2M}+\frac{L\eta^{2}}{2}{\left\lVert\frac{1}{M}\sum_{m=1}^{M}\nabla f(u_{m,r,h})\right\rVert}^{2}$
$\begin{split}&\leq f(u_{r,h})-\frac{\eta}{2M}\sum_{m=1}^{M}\left[{\left\lVert\nabla f(u_{r,h})\right\rVert}^{2}+{\left\lVert\nabla f(u_{m,r,h})\right\rVert}^{2}-{\left\lVert\nabla f(u_{r,h})-\nabla f(u_{m,r,h})\right\rVert}^{2}\right]+\frac{L\eta^{2}\sigma^{2}}{2M}\\ &\qquad+\frac{L\eta^{2}}{2M}\sum_{m=1}^{M}{\left\lVert\nabla f(u_{m,r,h})\right\rVert}^{2}\end{split}$
$\begin{split}&=f(u_{r,h})-\frac{\eta}{2}{\left\lVert\nabla f(u_{r,h})\right\rVert}^{2}-\frac{\eta(1-L\eta)}{2M}\sum_{m=1}^{M}{\left\lVert\nabla f(u_{m,r,h})\right\rVert}^{2}+\frac{\eta}{2}V_{r,h}+\frac{L\eta^{2}\sigma^{2}}{2M},\end{split}$
where $V_{r,h}=\frac{1}{M}\sum_{m=1}^{M}{\left\lVert\nabla f(u_{r,h})-\nabla f(u_{m,r,h})\right\rVert}^{2}\leq\frac{L^{2}}{M}\sum_{m=1}^{M}{\left\lVert u_{r,h}-u_{m,r,h}\right\rVert}^{2}$. Taking unconditional expectation, dropping the ${\left\lVert\nabla f(u_{r,h})\right\rVert}^{2}$ term and using Lemma B.7 we have
$\mathbb{E}\left[f(u_{r,h+1})\right]\leq\mathbb{E}\left[f(u_{r,h})\right]-\frac{\eta(1-L\eta)}{2M}\sum_{m=1}^{M}\mathbb{E}\left[\left\lVert\nabla f(u_{m,r,h})\right\rVert^{2}\right]+\frac{L\eta^{2}\sigma^{2}}{2M}+\eta^{3}L^{2}\sigma^{2}h.$
Observe that in the current scheme, $\bar{g}_{m,r,h}=\nabla f(u_{m,r,h})$. Suppose that $1-L\eta\geq\frac{1}{2}$, using this and telescoping yields
$\mathbb{E}\left[f(u_{r+1})\right]\leq\mathbb{E}\left[f(u_{r,h})\right]-\frac{\eta}{4M}\sum_{m=1}^{M}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}\left[\left\lVert\bar{g}_{m,r,h^{\prime}}\right\rVert^{2}\right]+\frac{L\eta^{2}\sigma^{2}(H-h)}{2M}+\eta^{3}L^{2}\sigma^{2}\sum_{h^{\prime}=h}^{H-1}h^{\prime}.$
Using Jensen’s inequality on $u_{r,h}=\frac{1}{M}\sum_{m=1}^{M}u_{m,r,h}=\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h}$ we obtain
$\mathbb{E}\left[f(u_{r+1})\right]\leq\frac{1}{M}\sum_{m=1}^{M}\left[f(y_{m,r,h})-\frac{\eta}{4}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}\left[\left\lVert\bar{g}_{m,r,h}\right\rVert^{2}\right]\right]+\frac{L\eta^{2}\sigma^{2}(H-h)}{2M}+L^{2}\eta^{3}\sigma^{2}\sum_{h^{\prime}=h}^{H-1}h^{\prime}.$
Summing up both sides as $h$ varies from 0 to H1H-1 we get
$Hf(u_{r+1})\leq\frac{1}{M}\sum_{m,h<H}\left[\mathbb{E}\left[f(y_{m,r,h})\right]-\frac{\eta}{4}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}\left[\left\lVert\bar{g}_{m,r,h}\right\rVert^{2}\right]\right]+\frac{L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{\eta^{3}L^{2}\sigma^{2}H^{3}}{2}.$
Define $G_{r}=\sum_{h=0}^{H-1}g_{r,h}$ and $\bar{G}_{r}=\sum_{h=0}^{H-1}\bar{g}_{r,h}$. The following lemma characterizes the evolution of the momentum sequence $z_{1},z_{2},\ldots$.
Lemma B.9 (Momentum sequence bound).
For any $r\geq 0$, the momentum sequence satisfies:
$\mathbb{E}_{r}\left[{\left\lVert z_{r+1}-x_{\ast}\right\rVert}^{2}\right]$ $={\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}+\gamma_{r}^{2}\eta^{2}\mathbb{E}_{r}\left[{\left\lVert\bar{G}_{r}\right\rVert}^{2}\right]+\frac{\gamma_{r}^{2}\eta^{2}\sigma^{2}}{M}-\gamma_{r}\eta\left\langle z_{r}-x_{\ast},\mathbb{E}_{r}\left[\bar{G}_{r}\right]\right\rangle.$
Proof.
Expanding the square,
${\left\lVert z_{r+1}-x_{\ast}\right\rVert}^{2}$ $={\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}+\gamma_{r}^{2}\eta^{2}{\left\lVert G_{r}\right\rVert}^{2}-2\gamma_{r}\eta\left\langle z_{r}-x_{\ast},G_{r}\right\rangle.$
Taking expectations and using Lemma B.3,
$\mathbb{E}_{r}\left[{\left\lVert z_{r+1}-x_{\ast}\right\rVert}^{2}\right]$ $={\left\lVert z_{r}-x_{\ast}\right\rVert}^{2}+\gamma_{r}^{2}\eta^{2}\mathbb{E}_{r}\left[{\left\lVert\bar{G}_{r}\right\rVert}^{2}\right]+\frac{\gamma_{r}^{2}\eta^{2}\sigma^{2}H}{M}-\gamma_{r}\eta\left\langle z_{r}-x_{\ast},\mathbb{E}_{r}\left[\bar{G}_{r}\right]\right\rangle.$
Proof of Theorem˜3.6.
Define the potential function
$\Phi_{r}=r(r+1)H(f(u_{r})-f(x_{*}))+\frac{2}{\gamma\eta}\parallel z_{r}-x_{*}\parallel^{2}.$
Using Lemma B.8 and Lemma B.9, we have
$\mathbb{E}_{r}[\Phi_{r+1}]-\Phi_{r}$
$=(r+1)(r+2)H\left(\mathbb{E}_{r}[f(u_{r+1})]-f(x_{*})\right)-r(r+1)H\left(f(u_{r})-f(x_{*})\right)$
$\qquad+\frac{2}{\gamma\eta}\left[\mathbb{E}_{r}[\parallel z_{r+1}-x_{*}\parallel^{2}]-\parallel z_{r}-x_{*}\parallel^{2}\right]$
$\leq(r+1)(r+2)\left[\frac{1}{M}\sum_{m,h<H}\left[(\mathbb{E}_{r}[f(y_{m,r,h})]-f(x_{*}))-\frac{\eta}{4}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}_{r}[\parallel\bar{g}_{m,r,h^{\prime}}\parallel^{2}]\right]+\frac{L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{\eta^{3}L^{2}\sigma^{2}H^{3}}{2}\right]$
$\qquad-r(r+1)H\left(f(u_{r})-f(x_{*})\right)$
$\qquad+\frac{\gamma\eta(r+1)^{2}}{2}\mathbb{E}_{r}[\parallel\bar{G}_{r}\parallel^{2}]+\frac{\gamma\eta\sigma^{2}(r+1)^{2}H}{2M}-2(r+1)\langle z_{r}-x_{*},\mathbb{E}_{r}[\bar{G}_{r}]\rangle$
$=\underbrace{\frac{1}{M}\sum_{m,h<H}\left[2(r+1)(\mathbb{E}_{r}[f(y_{m,r,h})]-f(x_{*}))+r(r+1)(\mathbb{E}_{r}[f(y_{m,r,h})]-f(u_{r}))-2(r+1)\langle z_{r}-x_{*},\mathbb{E}_{r}[\bar{g}_{m,r,h}]\rangle\right]}_{=:A}$
$\qquad\underbrace{-\frac{(r+1)(r+2)\eta}{4M}\sum_{m,h<H}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}_{r}[\parallel\bar{g}_{m,r,h^{\prime}}\parallel^{2}]+\frac{\gamma\eta(r+1)^{2}}{2}\mathbb{E}_{r}[\parallel\bar{G}_{r}\parallel^{2}]}_{=:B}$
$\qquad+\frac{(r+1)(r+2)L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{(r+1)(r+2)\eta^{3}L^{2}\sigma^{2}H^{3}}{2}+\frac{\gamma\eta\sigma^{2}(r+1)^{2}H}{2M}.$
Now, we bound the terms above separately. First, we bound $A$. Fix any $m,h<H$. We have, using convexity of $f$,
$2(r+1)(f(y_{m,r,h})-f(x_{*}))+r(r+1)(f(y_{m,r,h})-f(u_{r}))-2(r+1)\langle z_{r}-x_{*},\bar{g}_{m,r,h}\rangle$
$\leq 2(r+1)\langle y_{m,r,h}-x_{*},\bar{g}_{m,r,h}\rangle+r(r+1)\langle y_{m,r,h}-u_{r},\bar{g}_{m,r,h}\rangle-2(r+1)\langle z_{r}-x_{*},\bar{g}_{m,r,h}\rangle$
$=\langle(r+1)(r+2)y_{m,r,h}-r(r+1)u_{r}-2(r+1)z_{r},\bar{g}_{m,r,h}\rangle$
$=(r+1)(r+2)\langle y_{m,r,h}-x_{r},\bar{g}_{m,r,h}\rangle$
$=-\eta(r+1)(r+2)\sum_{h^{\prime}<h}\langle g_{m,r,h^{\prime}},\bar{g}_{m,r,h}\rangle.$
Hence,
$2(r+1)(\mathbb{E}_{r}[f(y_{m,r,h})]-f(x_{*}))+r(r+1)(\mathbb{E}_{r}[f(y_{m,r,h})]-f(u_{r}))-2(r+1)\langle z_{r}-x_{*},\mathbb{E}_{r}[\bar{g}_{m,r,h}]\rangle$
$\leq-\eta(r+1)(r+2)\sum_{h^{\prime}<h}\mathbb{E}_{r}[\langle\bar{g}_{m,r,h^{\prime}},\bar{g}_{m,r,h}\rangle]$
Since $A$ equals the sum of the above over all $m,h<H$ and dividing by $M$, we get:
$A$ $=\frac{-\eta(r+1)(r+2)}{M}\sum_{m,h<H}\sum_{h^{\prime}<h}\mathbb{E}_{r}[\langle\bar{g}_{m,r,h^{\prime}},\bar{g}_{m,r,h}\rangle]$
$=\frac{-\eta(r+1)(r+2)}{2M}\sum_{m}\mathbb{E}_{r}\left[\parallel\sum_{h<H}\bar{g}_{m,r,h}\parallel^{2}-\sum_{h<H}\parallel\bar{g}_{m,r,h}\parallel^{2}\right],$
where in the last line we used the algebraic identity that for any sequence of vectors $v_{0},\ldots,v_{H-1}$,
$\sum_{h<H}\sum_{s<h}\langle v_{s},v_{h}\rangle=\frac{1}{2}\left[\parallel\sum_{h<H}v_{h}\parallel^{2}-\sum_{h<H}\parallel v_{h}\parallel^{2}\right].$
Next, we have
$B$ $=-\frac{(r+1)(r+2)\eta}{4M}\sum_{m,h<H}\sum_{h^{\prime}=h}^{H-1}\mathbb{E}_{r}[\parallel\bar{g}_{m,r,h^{\prime}}\parallel^{2}]+\frac{\gamma\eta(r+1)^{2}}{2}\mathbb{E}_{r}[\parallel\bar{G}_{r}\parallel^{2}]$
$\leq-\frac{(r+1)(r+2)\eta}{4M}\sum_{m}\sum_{h<H}\mathbb{E}_{r}[\parallel\bar{g}_{m,r,h}\parallel^{2}]+\frac{\gamma\eta(r+1)^{2}}{2M}\sum_{m}\mathbb{E}_{r}\left[\parallel\sum_{h<H}\bar{g}_{m,r,h}\parallel^{2}\right].$
Hence, we have
$A+B$ $\leq\frac{-\eta(r+1)(r+2)}{2M}\sum_{m}\mathbb{E}_{r}\left[\parallel\sum_{h<H}\bar{g}_{m,r,h}\parallel^{2}-\sum_{h<H}\parallel\bar{g}_{m,r,h}\parallel^{2}\right]$
$\qquad-\frac{(r+1)(r+2)\eta}{4M}\sum_{m}\sum_{h<H}\mathbb{E}_{r}[\parallel\bar{g}_{m,r,h}\parallel^{2}]+\frac{\gamma\eta(r+1)^{2}}{2M}\sum_{m}\mathbb{E}_{r}\left[\parallel\sum_{h<H}\bar{g}_{m,r,h}\parallel^{2}\right]$
[…]
since $\gamma\leq 1$ implies $\gamma(r+1)-(r+2)=(r+1)(\gamma-1)-1\leq-1<0$, and the second term has a positive coefficient with a negative sign.
So overall, we have
$\mathbb{E}_{r}[\Phi_{r+1}]-\Phi_{r}$ $\leq\frac{(r+1)(r+2)L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{(r+1)(r+2)\eta^{3}L^{2}\sigma^{2}H^{3}}{2}+\frac{\gamma\eta\sigma^{2}(r+1)^{2}H}{2M}$
$\leq\frac{R^{2}L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{R^{2}\eta^{3}L^{2}\sigma^{2}H^{3}}{2}+\frac{\gamma\eta\sigma^{2}R^{2}H}{2M}.$
Summing up from $r=0$ to R1R-1, and taking expectations, we get
$\mathbb{E}[\Phi_{R}]-\Phi_{0}\leq\frac{R^{3}L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{R^{3}\eta^{3}L^{2}\sigma^{2}H^{3}}{2}+\frac{\gamma\eta\sigma^{2}R^{3}H}{2M}.$
Thus,
$R^{2}H(\mathbb{E}[f(u_{R})]-f(x_{*}))$
$\qquad\leq\mathbb{E}[\Phi_{R}]\leq\frac{2\parallel x_{0}-x_{*}\parallel^{2}}{\gamma\eta}+\frac{R^{3}L\eta^{2}\sigma^{2}H^{2}}{2M}+\frac{R^{3}\eta^{3}L^{2}\sigma^{2}H^{3}}{2}+\frac{\gamma\eta\sigma^{2}R^{3}H}{2M},$
which implies that
$\mathbb{E}[f(u_{R})]-f(x_{*})\leq\frac{2\parallel x_{0}-x_{*}\parallel^{2}}{\gamma\eta R^{2}H}+\frac{RL\eta^{2}\sigma^{2}H}{2M}+\frac{RL^{2}\eta^{3}\sigma^{2}H^{2}}{2}+\frac{\gamma\eta\sigma^{2}R}{2M}.$
The proof of Corollary˜3.7 is straightforward by substitution and is omitted for brevity.
B.5 Data-dependent guarantees
Lemma B.10.
Let $f$ be a convex and $L$-smooth function. Suppose that we run SGD on $f$ on $M$ parallel nodes as follows
$y_{m,r,0}$ $=x_{r},$
$y_{m,r,h+1}$ $=y_{m,r,h}-\eta g_{m,r,h},$
where $m=1,2,\ldots,M$, $h=0,1,\ldots,H-1$, and $g_{1,r,h},g_{2,r,h},\ldots,g_{M,r,h}$ are i.i.d. stochastic gradient estimates such that $\mathbb{E}_{r,h}\left[g_{m,r,h}\right]=\nabla f(y_{m,r,h})$, where $\mathbb{E}_{r,h}\left[\cdot\right]$ denotes expectation conditional on all information up to and including round $r$ and local step $h$, and $\left\lVert g_{m,r,h}-\nabla f(y_{m,r,h})\right\rVert\leq\sigma$. Define further $y_{r,h}=\frac{1}{M}\sum_{m=1}^{M}y_{m,r,h}$. Let $V_{r,h}=\frac{1}{M}\sum_{m=1}^{M}{\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}^{2}$. Then for all $\eta\leq\frac{1}{L}$ we have with probability at least $1-\delta$ that for all $h=0,1,\ldots,H$
$V_{r,h}\leq 4104\eta^{2}\sigma^{2}(h+1)\theta^{2}_{h-1,\delta},$
where $\theta_{h,\delta}=\log\frac{60\log 6h}{\delta}$.
Proof.
Define
$\Lambda_{r,h+1}$ $=\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}{\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}.$ (33)
We will bound $\Lambda_{r,h}$ first, and then use it to bound $V_{r,h}$ later. We have
$y_{m,r,h+1}-y_{s,r,h+1}=y_{m,r,h}-\eta g_{m,r,h}-\left[y_{s,r,h}-\eta g_{s,r,h}\right]$
$\quad=y_{m,r,h}-\eta\nabla f(y_{m,r,h})-\eta\left[g_{m,r,h}-\nabla f(y_{m,r,h})\right]-\left[y_{s,r,h}-\eta\nabla f(y_{s,r,h})-\eta\left[g_{s,r,h}-\nabla f(y_{s,r,h})\right]\right]$
$\quad=\left[y_{m,r,h}-\eta\nabla f(y_{m,r,h})-\left[y_{s,r,h}-\eta\nabla f(y_{s,r,h})\right]\right]-\eta\left[\left(g_{m,r,h}-g_{s,r,h}\right)-\left[\nabla f(y_{m,r,h})-\nabla f(y_{s,r,h})\right]\right].$
Therefore
$\begin{split}&{\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}=\quad\parallel T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})\parallel^{2}\\ &\qquad+\eta^{2}\parallel(g_{m,r,h}-g_{s,r,h})-(\nabla f(y_{m,r,h})-\nabla f(y_{s,r,h}))\parallel^{2}\\ &\qquad-2\eta\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),(g_{m,r,h}-g_{s,r,h})-(\nabla f(y_{m,r,h})-\nabla f(y_{s,r,h}))\right\rangle\end{split}$ (34)
We define $\rho_{m,r,h}$ as the stochastic gradient noise on node $m$ at round $r$, step $h$: $\rho_{m,r,h}=g_{m,r,h}-\nabla f(y_{m,r,h})$. Then we can write Equation˜34 as
${\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}$ $=\quad\parallel T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})\parallel^{2}+\eta^{2}\parallel\rho_{m,r,h}-\rho_{s,r,h}\parallel^{2}$
$\qquad-2\eta\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$ (35)
We now use the inequality ${\left\lVert a+b\right\rVert}^{2}\leq 2{\left\lVert a\right\rVert}^{2}+2{\left\lVert b\right\rVert}^{2}$ to get
$\begin{split}{\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}&\leq\parallel T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h})\parallel^{2}+2\eta^{2}\parallel\rho_{m,r,h}\parallel^{2}+2\eta^{2}{\left\lVert\rho_{s,r,h}\right\rVert}^{2}\\ &\qquad-2\eta\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.\end{split}$
By Lemma B.1, we have
${\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}$ $\leq{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}+2\eta^{2}{\left\lVert\rho_{m,r,h}\right\rVert}^{2}+2\eta^{2}{\left\lVert\rho_{s,r,h}\right\rVert}^{2}$
$\quad-2\eta\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$
Now, we consider the inner product term, observe
$\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle$
$\qquad=\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h})+T_{\eta}(y_{r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle$
$\qquad=\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle+\left\langle T_{\eta}(y_{r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle$
$\qquad=\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle+\left\langle-(T_{\eta}(y_{s,r,h})-T_{\eta}(y_{r,h})),-(\rho_{s,r,h}-\rho_{m,r,h})\right\rangle$
$\qquad=\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle+\left\langle T_{\eta}(y_{s,r,h})-T_{\eta}(y_{r,h}),\rho_{s,r,h}-\rho_{m,r,h}\right\rangle.$
Averaging with respect to $s$ and $m$
$\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}$ $\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h})+T_{\eta}(y_{r,h})-T_{\eta}(y_{s,r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle$
$\qquad=\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle$
$\qquad\qquad+\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{s,r,h})-T_{\eta}(y_{r,h}),\rho_{s,r,h}-\rho_{m,r,h}\right\rangle$
$\qquad=\frac{2}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$ (36)
Averaging Equation˜35 with respect to $m$ and $s$ and using Equation˜36 we get
$\begin{split}\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}{\left\lVert y_{m,r,h+1}-y_{s,r,h+1}\right\rVert}^{2}&\leq\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}+\frac{4\eta^{2}}{M}\sum_{m=1}^{M}{\left\lVert\rho_{m,r,h}\right\rVert}^{2}\\ &\qquad-\frac{2\eta}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.\end{split}$
Using $\Lambda_{r,h}$ as defined in Equation˜33 we obtain the recursion
$\Lambda_{r,h+1}$ $\leq\Lambda_{r,h}+\frac{4\eta^{2}}{M}\sum_{m=1}^{M}{\left\lVert\rho_{m,r,h}\right\rVert}^{2}-\frac{2\eta}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$
Now observe that ${\left\lVert\rho_{m,r,h}\right\rVert}^{2}\leq\sigma^{2}$ by assumption, therefore
$\Lambda_{r,h+1}$ $\leq\Lambda_{r,h}+4\eta^{2}\sigma^{2}-\frac{2\eta}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$
Recursing the above inequality we get
$\Lambda_{r,h}$ $\leq\Lambda_{r,0}+4\eta^{2}\sigma^{2}h-\frac{2\eta}{M^{2}}\sum_{k=0}^{h-1}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,k})-T_{\eta}(y_{r,k}),\rho_{m,r,k}-\rho_{s,r,k}\right\rangle$
$=4\eta^{2}\sigma^{2}h-\frac{2\eta}{M^{2}}\sum_{k=0}^{h-1}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,k})-T_{\eta}(y_{r,k}),\rho_{m,r,k}-\rho_{s,r,k}\right\rangle,$ (37)
where we used the fact that since $y_{m,r,0}=y_{s,r,0}=x_{r}$ for all $m,s$ then $\Lambda_{r,0}=0$. Define
$\mu_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,h}-y_{r,h}\right\rVert,$ $\overline{\mu}_{r,h}=\max_{k\leq h}\mu_{r,k},$ (38)
$X_{r,h}$ $=\frac{1}{\overline{\mu}_{r,h}}\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle.$ (39)
Let $\mathbb{E}_{r,h}\left[\cdot\right]$ denote the expectation conditional on all information up to and including round $r$ and local step $h$. Then,
$\mathbb{E}_{r,h}\left[X_{r,h}\right]$ $=0.$
Furthermore, we have by the triangle inequality, then our assumption on the noise followed by Lemma B.1 that almost surely
$\left\lvert\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle\right\rvert$ $\leq\left\lVert T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h})\right\rVert\left\lVert\rho_{m,r,h}-\rho_{s,r,h}\right\rVert$
$\leq\left\lVert T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h})\right\rVert\left(\left\lVert\rho_{m,r,h}\right\rVert+\left\lVert\rho_{s,r,h}\right\rVert\right)$
$\leq 2\sigma\left\lVert T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h})\right\rVert$
$\leq 2\sigma\left\lVert y_{m,r,h}-y_{r,h}\right\rVert.$ (40)
By the definition of $X_{r,h}$ (Equation˜39), the triangle inequality, Equation˜40, and the definition of $\overline{\mu}_{r,h}$ (Equation˜38) we have almost surely
$\left\lvert X_{r,h}\right\rvert$ $=\frac{1}{\overline{\mu}_{r,h}}\left\lvert\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle\right\rvert$
$\leq\frac{1}{\overline{\mu}_{r,h}}\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\lvert\left\langle T_{\eta}(y_{m,r,h})-T_{\eta}(y_{r,h}),\rho_{m,r,h}-\rho_{s,r,h}\right\rangle\right\rvert$
$\leq\frac{2\sigma}{\overline{\mu}_{r,h}}\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\lVert y_{m,r,h}-y_{r,h}\right\rVert$
$=2\sigma\frac{\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}{\overline{\mu}_{r,h}}$
$\leq 2\sigma.$
Then by Lemma˜B.4 with $y_{h}=\overline{\mu}_{r,h}$ we have with probability at least $1-\delta$
$\left\lvert\sum_{k=0}^{h-1}\overline{\mu}_{r,k}X_{r,k}\right\rvert$ $\leq 8\overline{\mu}_{r,h-1}\sqrt{\theta_{h-1,\delta}\sum_{k=0}^{h-1}X_{r,k}^{2}+4\sigma^{2}\theta_{h,\delta}^{2}}$
$\leq 8\overline{\mu}_{r,h-1}\sqrt{\theta_{h-1,\delta}4h\sigma^{2}+4\sigma^{2}\theta_{h,\delta}^{2}}$
$\leq 16\overline{\mu}_{r,h-1}\theta_{h-1,\delta}\sigma\sqrt{h+1}.$ (41)
Observe that
$\sum_{k=0}^{h-1}\overline{\mu}_{r,k}X_{r,k}$ $=\frac{1}{M^{2}}\sum_{k=0}^{h-1}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\langle T_{\eta}(y_{m,r,k})-T_{\eta}(y_{r,k}),\rho_{m,r,k}-\rho_{s,r,k}\right\rangle.$
Using this and Equation˜41 to upper bound the right hand side of Equation˜37 we obtain
$\Lambda_{r,h}$ $\leq 4\eta^{2}\sigma^{2}h+32\eta\overline{\mu}_{r,h-1}\theta_{h-1,\delta}\sigma\sqrt{h+1}$
$\leq 4\eta^{2}\sigma^{2}h+2\alpha(32\eta\theta_{h-1,\delta}\sigma\sqrt{h+1})^{2}+\frac{\overline{\mu}_{r,h-1}^{2}}{2\alpha}$
$=\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}(4+2048\alpha)+\frac{\overline{\mu}_{r,h-1}^{2}}{2\alpha},$ (42)
where we used that $2ab\leq\alpha a^{2}+\frac{1}{\alpha}b^{2}$ in the second step. Let $\overline{\Lambda}_{r,h}=\max_{k\leq h}\Lambda_{r,k}$. Observe that the right hand side of Equation˜42 is increasing in $h$, therefore
$\overline{\Lambda}_{r,h}\leq\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}(4+2048\alpha)+\frac{\overline{\mu}_{r,h-1}^{2}}{2\alpha}.$ (43)
Observe that by the triangle inequality followed by Lemma˜B.2
$\mu_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,h}-y_{r,h}\right\rVert$
$\leq\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert$
$\leq\sqrt{\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert^{2}}$
$=\sqrt{\Lambda_{r,h}}.$
It follows that $\overline{\mu}_{r,h}\leq\sqrt{\overline{\Lambda}_{r,h}}$. Using this in Equation˜43 we get
$\overline{\Lambda}_{r,h}$ $\leq\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}(4+2048\alpha)+\frac{\overline{\Lambda}_{r,h-1}}{2\alpha}$
$\leq\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}(4+2048\alpha)+\frac{\overline{\Lambda}_{r,h}}{2\alpha}.$
Rearranging we get
$\left(1-\frac{1}{2\alpha}\right)\overline{\Lambda}_{r,h}\leq\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}(4+2048\alpha)$
Put $\alpha=1$, then
$\overline{\Lambda}_{r,h}\leq 4104\eta^{2}\sigma^{2}(h+1)\theta_{h-1,\delta}^{2}.$ (44)
Now that we have our bound on $\overline{\Lambda}_{r,h}$, we can use it to bound $V_{r,h}$ as follows
$V_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}{\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}^{2}.$ (45)
Observe that by Jensen’s inequality
${\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}^{2}$ $={\left\lVert y_{m,r,h}-\frac{1}{M}\sum_{s=1}^{M}y_{s,r,h}\right\rVert}^{2}$
$={\left\lVert\frac{1}{M}(y_{m,r,h}-y_{s,r,h})\right\rVert}^{2}$
$\leq\frac{1}{M}\sum_{s=1}^{M}{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}.$ (46)
Combining Equations˜45 and 46 we have
$V_{r,h}$ $\leq\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{s=1}^{M}{\left\lVert y_{m,r,h}-y_{s,r,h}\right\rVert}^{2}=\Lambda_{r,h}.$
Combining this with Equation˜44 yields the lemma’s statement. ∎
Lemma B.11.
(Per-round regret). In Algorithm 1, the iterates in a single communication round satisfy
$\begin{split}{\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\eta\left\lvert 1-\gamma\right\rvert\zeta_{2}\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\\ &\qquad+\frac{\gamma\zeta_{3}H}{\alpha}+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle,\end{split}$
where $\alpha>0$ is arbitrary and
$\zeta_{2}=\max_{h}\left\lVert y_{r,h}-y_{r,0}\right\rVert,$ $\zeta_{3}=\max_{h}\frac{1}{M}\sum_{m=1}^{M}{\left\lVert y_{m,r,h}-y_{r,h}\right\rVert}^{2}.$
Proof.
Define the virtual sequences
$g_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}g_{m,r,h},$ $x_{r,0}=x_{r},$ $x_{r,h+1}=x_{r,h}-\gamma\eta g_{r,h}.$
We have
${\left\lVert x_{r,h+1}-x_{\ast}\right\rVert}^{2}$ $={\left\lVert x_{r,h}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}{\left\lVert g_{r,h}\right\rVert}^{2}-2\gamma\eta\left\langle x_{r,h}-x_{\ast},g_{r,h}\right\rangle$ (47)
The inner product term can be decomposed as
$-\left\langle x_{r,h}-x_{\ast},g_{r,h}\right\rangle$ $=-\left\langle x_{r,h}-y_{r,h},g_{r,h}\right\rangle-\left\langle y_{r,h}-x_{\ast},g_{r,h}\right\rangle.$ (48)
Observe that $x_{r,h}=x_{r}-\gamma\eta\sum_{s=0}^{h-1}g_{r,s}$ and $y_{r,h}=x_{r}-\eta\sum_{s=0}^{h-1}g_{r,s}$. Therefore,
$\left\lVert x_{r,h}-y_{r,h}\right\rVert$ $=\left\lVert(\gamma-1)\eta\sum_{s=0}^{h-1}g_{r,s}\right\rVert$
$=\left\lvert\gamma-1\right\rvert\left\lVert y_{r,h}-y_{r,0}\right\rVert$
$\leq\left\lvert\gamma-1\right\rvert\zeta_{2},$
where $\zeta_{2}=\max_{h}\left\lVert y_{r,h}-y_{r,0}\right\rVert$. Using this in Equation˜48
$-\left\langle x_{r,h}-y_{r,h},g_{r,h}\right\rangle\leq\left\lVert x_{r,h}-y_{r,h}\right\rVert\left\lVert g_{r,h}\right\rVert\leq\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert.$ (49)
Plugging Equation˜49 into Equation˜48 we get
- $\left\langle x_{r,h}-x_{\ast},g_{r,h}\right\rangle\leq\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert-\left\langle y_{r,h}-x_{\ast},g_{r,h}\right\rangle$
$\quad=\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert-\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{r,h}-x_{\ast},g_{m,r,h}\right\rangle$
$\quad=\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert-\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{r,h}-y_{m,r,h},g_{m,r,h}\right\rangle-\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.$ (50)
For the second term in Equation˜50 we have
$-\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{r,h}-y_{m,r,h},g_{m,r,h}\right\rangle$ $\leq\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{r,h}-y_{m,r,h}\right\rVert\left\lVert g_{m,r,h}\right\rVert$
$\leq\frac{1}{M}\sum_{m=1}^{M}\left[\frac{{\left\lVert y_{r,h}-y_{m,r,h}\right\rVert}^{2}}{2\alpha\eta}+\frac{\alpha\eta}{2}{\left\lVert g_{m,r,h}\right\rVert}^{2}\right]$
$\leq\frac{\zeta_{3}}{2\alpha\eta}+\frac{\alpha\eta}{2}\frac{1}{M}\sum_{m=1}^{M}{\left\lVert g_{m,r,h}\right\rVert}^{2}.$ (51)
Plugging Equation˜51 into Equation˜50 we get
$\begin{split}-\left\langle x_{r,h}-x_{\ast},g_{r,h}\right\rangle&\leq\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert+\frac{\zeta_{3}}{2\alpha\eta}+\frac{\alpha\eta}{2}\frac{1}{M}\sum_{m=1}^{M}{\left\lVert g_{m,r,h}\right\rVert}^{2}\\ &\quad-\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.\end{split}$ (52)
Plug Equation˜52 back into Equation˜47 to get
$\begin{split}{\left\lVert x_{r,h+1}-x_{\ast}\right\rVert}^{2}&\leq{\left\lVert x_{r,h}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\eta\left\lvert 1-\gamma\right\rvert\zeta_{2}\left\lVert g_{r,h}\right\rVert\\ &\qquad+\frac{\gamma\zeta_{3}}{\alpha}+\frac{\alpha\gamma\eta^{2}}{M}\sum_{m=1}^{M}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.\end{split}$
Recursing we get
$\begin{split}{\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\eta\left\lvert 1-\gamma\right\rvert\zeta_{2}\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\\ &\qquad+\frac{\gamma\zeta_{3}H}{\alpha}+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.\end{split}$
Proof of Theorem˜3.8.
Starting with the per-round recursion lemma, we have
$\begin{split}{\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\eta\left\lvert 1-\gamma\right\rvert\zeta_{2}\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\\ &\qquad+\frac{\gamma\zeta_{3}H}{\alpha}+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.\end{split}$
Observe that
$\left\lVert y_{r,h}-y_{r,0}\right\rVert$ $=\eta\left\lVert\sum_{k=0}^{h-1}g_{r,k}\right\rVert$
$\leq\eta\sum_{k=0}^{h-1}\left\lVert g_{r,k}\right\rVert$
$\leq\eta\sum_{k=0}^{H-1}\left\lVert g_{r,k}\right\rVert.$ (53)
Since this holds for any $h$, we have that $\zeta_{2}\leq\eta\sum_{k=0}^{H-1}\parallel g_{r,k}\parallel$, where $\zeta_{2}$ is defined in Lemma˜B.11. Moreover, by Lemma˜B.10 we have that with probability $1-\delta$ and an application of the union bound that for all $r,h$
$\frac{1}{M}\sum_{m=1}^{M}\parallel y_{m,r,h}-y_{r,h}\parallel^{2}\leq 4104\iota\eta^{2}\sigma^{2}H,$ (54)
where $\iota=2\cdot\log\frac{60\log 6RH}{\delta}$ and we used that $H+1\leq 2H$. Since this bound holds for all $h$, we have
$\zeta_{3}=\max_{h}\frac{1}{M}\sum_{m=1}^{M}\parallel y_{m,r,h}-y_{r,h}\parallel^{2}\leq 4104\iota\eta^{2}\sigma^{2}H.$
Therefore by Equation˜53 and Lemma˜B.10
${\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\left\lvert 1-\gamma\right\rvert\eta^{2}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}$
$\qquad+\frac{4104\gamma\eta^{2}\sigma^{2}H^{2}}{\alpha}\iota+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},g_{m,r,h}\right\rangle.$
Let $\xi_{m,r,h}=g_{m,r,h}-\nabla f(y_{m,r,h})$. Then,
$\begin{split}{\left\lVert x_{r+1}-x_{\ast}\right\rVert}^{2}&\leq{\left\lVert x_{r}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\left\lvert 1-\gamma\right\rvert\eta^{2}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}+\frac{4104\gamma\eta^{2}\sigma^{2}H^{2}}{\alpha}\iota\\ &\qquad+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}{\left\lVert g_{m,r,h}\right\rVert}^{2}-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\\ &\qquad-\frac{2\gamma\eta}{M}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\xi_{m,r,h}\right\rangle,\end{split}$ (55)
where $\xi_{m,r,h}=g_{m,r,h}-\nabla f(y_{m,r,h})$. Define
$\nu_{r,h}$ $=\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,h}-x_{\ast}\right\rVert,$ $\overline{\nu}_{r,h}=\max_{p\leq r,s\leq h}\nu_{p,s}.$
Let
$X_{r,h}=\frac{1}{\bar{\nu}_{r,h}}\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\xi_{m,r,h}\right\rangle$
Let $\mathcal{F}_{r,h-1}$ denote the sigma algebra generated by all randomness up to and including step r,h1r,h-1. Note that
$\mathbb{E}_{\mathcal{F}_{r,h-1}}\left[X_{r,h}\right]$ $=\frac{1}{\bar{\nu}_{r,h}}\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{\mathcal{F}_{r,h}}\left[\left\langle y_{m,r,h}-x_{\ast},\xi_{m,r,h}\right\rangle\right]$
$=\frac{1}{\bar{\nu}_{r,h}}\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\mathbb{E}_{\mathcal{F}_{r,h}}\left[\xi_{m,r,h}\right]\right\rangle$
$=0,$
where we used that $\nu_{r,h}$ and $y_{m,r,h}$ are both $\mathcal{F}_{r,h-1}$-measurable and that the noise has mean zero. The edge cases $X_{r,0}$ are handled similarly. Moreover, using the assumption that $\left\lVert\xi_{m,r,h}\right\rVert\leq\sigma$ almost surely and the definition of $\bar{\nu}_{r,h}$,
$\left\lVert X_{r,h}\right\rVert$ $=\left\lVert\frac{1}{\bar{\nu}_{r,h}}\frac{1}{M}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\xi_{m,r,h}\right\rangle\right\rVert$
$\leq\frac{1}{M}\sum_{m=1}^{M}\frac{\left\lVert y_{m,r,h}-x_{\ast}\right\rVert\left\lVert\xi_{m,r,h}\right\rVert}{\bar{\nu}_{r,h}}$
$\leq\frac{1}{M}\sum_{m=1}^{M}(1\cdot\sigma)$
$=\sigma.$
Applying Lemma˜B.4 on $X_{r,h}$ with $y_{r,h}=\bar{\nu}_{r,h}$, $C_{r,h}=\sigma$, $\hat{X}_{r,h}=0$ we have
$\left\lvert\frac{1}{M}\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}\sum_{m=1}^{M}\left\langle y_{m,r,h}-x_{\ast},\xi_{m,r,h}\right\rangle\right\rvert\leq 16\overline{\nu}_{R,H}\iota\sigma\sqrt{RH},$ (56)
where $\iota$ is defined as before. Using Equation˜56 in Equation˜55
$\begin{split}&\frac{2\gamma\eta}{M}\sum_{m,r,h}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\leq{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}-{\left\lVert x_{R}-x_{\ast}\right\rVert}^{2}+\gamma^{2}\eta^{2}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}\\ &\qquad+2\gamma\left\lvert 1-\gamma\right\rvert\eta^{2}\sum_{r=0}^{R-1}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}+R\cdot\frac{4104\gamma\eta^{2}\sigma^{2}H^{2}}{\alpha}\iota\\ &\qquad+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}+2\gamma\eta\left[16\overline{\nu}_{R,H}\iota\sigma\sqrt{RH}\right].\end{split}$ (57)
Let
$\begin{split}\Omega&=\gamma^{2}\eta^{2}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}+2\gamma\left\lvert 1-\gamma\right\rvert\eta^{2}\sum_{r=0}^{R-1}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}+R\cdot\frac{4104\gamma\eta^{2}\sigma^{2}H^{2}}{\alpha}\iota\\ &+\frac{\alpha\gamma\eta^{2}}{2}\frac{1}{M}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}\end{split}$ (58)
Then by convexity and Equation˜57 we get
${\left\lVert x_{R}-x_{\ast}\right\rVert}^{2}$ $\leq{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}+\Omega+2\gamma\eta\left[16\overline{\nu}_{R,H}\iota\sigma\sqrt{RH}\right]-\frac{2\gamma\eta}{M}\sum_{m,r,h}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle$
$\leq{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}+\Omega+2\gamma\eta\left[16\overline{\nu}_{R,H}\iota\sigma\sqrt{RH}\right],$ (59)
where in the second line we used that $x_{\ast}$ is the minimizer of $f$ and therefore $\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\geq 0$ by convexity. It is not difficult to see that this guarantee in fact applies not just on ${\left\lVert x_{R}-x_{\ast}\right\rVert}^{2}$ but on any $x_{r}$. Let $d_{r}=\left\lVert x_{r}-x_{\ast}\right\rVert$ and $\overline{d}_{r}=\max_{r^{\prime}\leq r}d_{r^{\prime}}$. Observe
$\nu_{r,h}=\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,h}-x_{\ast}\right\rVert$ $\leq\frac{1}{M}\sum_{m=1}^{M}\left[\left\lVert y_{m,r,h}-y_{m,r,0}\right\rVert+\left\lVert x_{r}-x_{\ast}\right\rVert\right]$
$\leq\left[\frac{\eta}{M}\sum_{m=1}^{M}\sum_{k=0}^{h-1}\left\lVert g_{m,r,k}\right\rVert\right]+\left\lVert x_{r}-x_{\ast}\right\rVert$
$\leq\left[\frac{\eta}{M}\sum_{m=1}^{M}\sum_{k=0}^{H-1}\left\lVert g_{m,r,k}\right\rVert\right]+\left\lVert x_{r}-x_{\ast}\right\rVert.$ (60)
Using Equation˜60 in Equation˜59 we get
$\overline{d}_{R}^{2}$ $\leq d_{0}^{2}+\Omega+32\gamma\eta\iota\sigma\sqrt{RH}\overline{\nu}_{R,H}$
$\leq d_{0}^{2}+\Omega+32\gamma\eta\iota\sigma\sqrt{RH}\left[\frac{\eta}{M}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right]+32\gamma\eta\iota\sigma\sqrt{RH}\overline{d}_{R}$
$\leq d_{0}^{2}+\Omega+2\left(32\gamma\eta\iota\sigma\sqrt{RH}\right)^{2}+\eta^{2}\left(\frac{1}{M}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}+\frac{\overline{d}_{R}^{2}}{2}.$
Therefore
$\overline{d}_{R}^{2}$ $\leq 2d_{0}^{2}+2\Omega+4096\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+2\eta^{2}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}.$ (61)
By the triangle inequality applied twice and the definition of $\bar{d}_{R}$,
$\left\lVert y_{m,r,s}-x_{\ast}\right\rVert$ $\leq\left\lVert y_{m,r,0}-y_{m,r,s}\right\rVert+\left\lVert y_{m,r,0}-x_{\ast}\right\rVert$
$=\eta\left\lVert\sum_{h=0}^{s-1}g_{m,r,h}\right\rVert+\left\lVert y_{m,r,0}-x_{\ast}\right\rVert$
$\leq\eta\sum_{h=0}^{s-1}\left\lVert g_{m,r,h}\right\rVert+\left\lVert y_{m,r,0}-x_{\ast}\right\rVert$
$\leq\eta\sum_{h=0}^{s-1}\left\lVert g_{m,r,h}\right\rVert+\bar{d}_{R}$
$\leq\eta\sum_{h=0}^{H-1}\left\lVert g_{m,r,h}\right\rVert+\bar{d}_{R}.$
Therefore
$\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,s}-x_{\ast}\right\rVert\leq\eta\left(\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\left\lVert g_{m,r,h}\right\rVert\right)+\bar{d}_{R}$
We now use the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ to get
$\nu_{r,s}^{2}$ $=\left(\frac{1}{M}\sum_{m=1}^{M}\left\lVert y_{m,r,s}-x_{\ast}\right\rVert\right)^{2}$
$\leq 2\left(\eta\left(\frac{1}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\left\lVert g_{m,r,h}\right\rVert\right)\right)^{2}+2\bar{d}_{R}^{2}$
$=2\eta^{2}\left(\frac{1}{M}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}+2\overline{d}_{R}^{2}.$
Finally, using our bound on $\bar{d}_{R}^{2}$ given by equation (61)
$\nu_{r,s}^{2}$ $\leq 4d_{0}^{2}+4\Omega+8192\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+6\eta^{2}\left(\frac{1}{M}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2},$
Therefore
$\bar{\nu}_{R,H}^{2}$ $=\max_{r,s}\nu_{r,s}^{2}$
$\leq 4d_{0}^{2}+4\Omega+8192\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+6\eta^{2}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}.$
By Equations˜57 and 58 and the last equation,
$\frac{2\gamma\eta}{M}\sum_{m,r,h}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\leq{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}-{\left\lVert x_{R}-x_{\ast}\right\rVert}^{2}+\Omega+2\gamma\eta\left[16\overline{\nu}_{R,H}\iota\sigma\sqrt{RH}\right]$
$\qquad\leq d_{0}^{2}-d_{R}^{2}+\Omega+\frac{(32\gamma\eta\iota\sigma\sqrt{RH})^{2}}{2}+4\left[d_{0}^{2}+\Omega+2048\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH\right]+6\eta^{2}R\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}$
$\qquad=d_{0}^{2}-d_{R}^{2}+\Omega+\frac{(32\gamma\eta\iota\sigma\sqrt{RH})^{2}}{2}+4\left[d_{0}^{2}+\Omega+2048\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH\right]+6\eta^{2}R\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}$
$\qquad\leq d_{0}^{2}-d_{R}^{2}+6\Omega+8704\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+4d_{0}^{2}+6\eta^{2}R\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}.$ (62)
Dropping the $d_{R}^{2}$ term, we get
$\frac{2\gamma\eta}{M}\sum_{m,r,h}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\leq 5d_{0}^{2}+6\Omega+8704\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+6\eta^{2}R\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}$
$\begin{split}&\qquad\leq 5d_{0}^{2}+6\gamma^{2}\eta^{2}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}+12\gamma\left\lvert 1-\gamma\right\rvert\eta^{2}\sum_{r=0}^{R-1}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}+RH\frac{24624\gamma\eta^{2}\sigma^{2}H\iota}{\alpha}\\ &\qquad+\frac{3\alpha\gamma\eta^{2}}{M}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}+8704\gamma^{2}\eta^{2}\iota^{2}\sigma^{2}RH+6\eta^{2}R\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}.\end{split}$
Dividing both sides by $2\gamma\eta RH$ gives
$\begin{split}&\frac{1}{MRH}\sum_{m,r,h}\left\langle y_{m,r,h}-x_{\ast},\nabla f(y_{m,r,h})\right\rangle\leq\frac{5d_{0}^{2}}{2\gamma\eta RH}+\frac{3\gamma\eta}{RH}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}\\ &\qquad+\frac{6\left\lvert 1-\gamma\right\rvert\eta}{RH}\sum_{r=0}^{R-1}\left(\sum_{h=0}^{H-1}\left\lVert g_{r,h}\right\rVert\right)^{2}+\frac{24624\eta\sigma^{2}H\iota}{\alpha}\\ &\qquad+\frac{3\alpha\eta}{MRH}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}+8704\gamma\eta\iota^{2}\sigma^{2}+\frac{6\eta}{\gamma H}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}.\end{split}$ (63)
Observe that by optimizing over $\alpha$ we have
$\frac{24624\eta\sigma^{2}H\iota}{\alpha}+\frac{3\alpha\eta}{MRH}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}$ $\leq 2\sqrt{\left(24624\eta\sigma^{2}H\iota\right)\left(\frac{3\eta}{MRH}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}\right)}$
$\leq 544\eta\sigma\iota\sqrt{\frac{1}{MR}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}}.$
Using this in Equation˜63 followed by convexity completes the proof. ∎